### v1

# Quadratic Gaussian Splatting for Efficient and Detailed Surface Reconstruction

Abstract
Recently, 3D Gaussian Splatting (3DGS) has attracted attention for its superior rendering quality and speed over Neural Radiance Fields (NeRF). To address 3DGS’s limitations in surface representation, 2D Gaussian Splatting (2DGS) introduced disks as scene primitives to model and reconstruct geometries from multi-view images, offering view-consistent geometry. However, the disk’s first-order linear approximation often leads to over-smoothed results. We propose Quadratic Gaussian Splatting (QGS), a novel method that replaces disks with quadric surfaces, enhancing geometric fitting. QGS defines Gaussian distributions in non-Euclidean space, allowing primitives to capture more complex textures. As a second-order surface approximation, QGS also renders spatial curvature to guide the normal consistency term, to effectively reduce over-smoothing. Moreover, QGS is a generalized version of 2DGS that achieves more accurate and detailed reconstructions, as verified by experiments on DTU and TNT, demonstrating its effectiveness in surpassing most state-of-the-art methods in geometry reconstruction. Our code will be released as open source. The project is available at https://quadraticgs.github.io/QGS
Figure 1: Quadratic Gaussian Splatting, a novel quadric-based approach for recovering high-precision real-world geometry from multi-view RGB images. In the Courthouse scene, our method achieves multi-view consistent normals, depth, and curvature, producing high-quality meshes and realistic images. Compared to 2DGS, QGS more accurately models the building’s contours.
1 Introduction
Surface reconstruction and realistic novel view synthesis (NVS) are critical tasks in computer graphics and vision. Their goal is to recover dense geometric structures from images captured from different viewpoints and render photorealistic images. Recently, 3D Gaussian Splatting (3DGS) has surpassed Neural Radiance Fileds (NeRF) based methods [2, 23, 24] in both rendering quality and speed by combining traditional splatting techniques with end-to-end optimization. Subsequently, GS-based methods have seen rapid advancements in dynamic reconstruction [13, 8], editing [6, 29], and large-scale scene reconstruction [20, 21].
However, vanilla splatting ignores the Gaussian contribution along the z-axis and uses approximation [39, 16]. As a result, it can render high-quality, multi-view consistent textures even at offsets from the scene surface, making it challenging to capture accurate geometry. In subsequent work [35, 19, 12, 4, 36], ray-splat intersections are computed in 3D to retrieve Gaussian weights, a key factor for accurate scene geometry recovery. In methods with Gaussian ellipsoids as primitives [35, 36], ray-splat intersections depend on the view direction, limiting normal consistency across views. In contrast, methods using Gaussian disks as primitives [12, 19] can ensure multi-view consistent geometry due to their surface-based nature, facilitating the incorporation of multi-view geometric consistency [4, 19]. However, disks, as a first-order fit of geometry, often result in overly smooth reconstructions, as shown in the Fig 2.
Earlier works [26, 40, 31] demonstrated that surfel elements can effectively represent complex scene geometry, such as atomic models. Building on this, 2DGS [12] combined Gaussian distributions with disk elements to achieve state-of-the-art (SOTA) reconstruction, inspiring our approach. In this paper, we propose a novel surface-based rendering method, which we call Quadratic Gaussian Splatting (QGS), for 3D scene reconstruction and novel view synthesis. This method enhances the geometric fitting of scene primitives through higher-order representation, enabling the extraction of accurate surface mesh from the quadric model.
Unlike previous splatting techniques operating in Euclidean space, such as 3DGS or 2DGS, our QGS defines the Gaussian distribution on a quadratic paraboloid, which can continuously transition between convex and concave forms, offering greater geometric fitting flexibility. The key technique in using paraboloids for splatting lies in establishing Gaussian distributions based on geodesic distances in non-Euclidean space, which allows the distribution’s energy to be concentrated on the surface, to effectively capture complex geometric textures. Compared to the first-order approximation in 2DGS, QGS provides a second-order fit to the scene geometry. As a result, QGS not only produces multi-view consistent normals and depth but also provides curvature information, which represents surface bending and helps guide single-view normal consistency supervision. Furthermore, we observe that the volumetric rendering order used by most GS methods causes popping artifacts, which impact both novel view synthesis and geometric reconstruction. To address this, we adopted the sorting criteria from StopThePop [25], restructuring it to accommodate quadric surface structures, thereby improving the centroid sorting used in 2DGS to better handle complex intersections of quadratic surfaces. Finally, QGS is able to extract high-quality and detail-rich mesh models, as shown in the Fig 1.
In summary, the main contributions of our work are as follows:
  • We propose QGS, a novel and efficient differentiable representation that uses quadric surfaces as scene primitives, offering stronger geometric fitting capabilities.
  • We propose the first Gaussian Splatting work to introduce geodesic distance to establish Gaussian distributions on surfaces, enabling primitives to fit more complex textures.
  • With stricter depth sorting in the higher-degree-of-freedom QGS, our method achieves SOTA geometric reconstruction and enhanced rendering quality.
2 Related work
2.1 Novel View Synthesis
Novel view synthesis has seen rapid development in recent years, especially since the introduction of NeRF [23]. NeRF encodes scene geometry and view-dependent appearance using a multi-layer perceptron (MLP) and optimizes it end-to-end through volumetric rendering to produce realistic images. Subsequent methods introduced various improvements. Mip-NeRF [1], Mip-NeRF 360 [2], and Zip-NeRF [3] addressed NeRF’s aliasing issues by introducing new sampling strategies. I-NGP [24] and DVGO [28] significantly accelerated NeRF’s training and rendering speeds by encoding scenes into feature grids instead of MLPs.
Recently, 3DGS [16] has gained attention for its more realistic rendering capabilities and real-time rendering speed, surpassing NeRF in many cases [34, 25, 22, 38]. Mip-Splatting [34] introduced a 3D smoothing filter to eliminate high-frequency artifacts during rendering. StopThePop [25] proposed per-tile and per-pixel resorting to resolve multi-view geometric inconsistencies. Scaffold-GS [22] effectively reduced redundant Gaussians and improved rendering quality by constructing anchor points to distribute local 3D Gaussians.
2.2 Neural Surface Reconstruction
Compared to traditional multi-view geometric reconstruction, neural volumetric rendering methods often produce smoother and more complete results. NeuS [30] and VolSDF [32] introduced signed distance fields (SDF) to describe scene geometry based on NeRF [23]. Geo-NeuS [9] and NeuralWarp [10] improved reconstruction by incorporating multi-view consistency. Neuralangelo [18] reduced MLP reliance by using feature grids and numerical gradients to improve geometric reconstruction quality.
However, neural rendering-based surface reconstruction generally requires several hours to converge using only images. Additionally, due to the structured implicit fields, accessing and editing scene geometry is not straightforward.
2.3 Gaussian Splatting Surface Reconstruction
With the rapid development of 3DGS across various fields, Gaussian splatting methods have also made significant progress in surface reconstruction. SuGaR [11] introduced regularization terms to encourage Gaussians to fit the scene surface by building volumetric density fields from Gaussian ellipsoids and using Poisson reconstruction [15] to obtain a mesh. GSDF [33] and NeuSG [5] combined 3D Gaussian ellipsoids with SDF, achieving high-quality geometric reconstructions and rendering results. Rade-GS [36], GOF [35], and PGSR [4] computed ray-Gaussian ellipsoid intersections to obtain unbiased depth and used the normal consistency supervision from 2DGS [12] to achieve state-of-the-art reconstruction results. 2DGS flattened Gaussian ellipsoids into Gaussian disks to better align Gaussian primitives with surfaces and introduced two regularization losses to provide additional geometric constraints. MVG-splatting [19] extended 2DGS by introducing multi-view consistency constraints, enabling more complex surface reconstructions. In this paper, we propose Quadratic Gaussian Splatting (QGS), a new surface representation based on quadric surfaces designed to improve the local geometric fitting capability of primitives. We further establish Gaussian distributions on quadric, enabling end-to-end optimization. Lastly, we address the volume rendering order issue in 2DGS by incorporating the sorting method from StopThePop [25], achieving precise geometric reconstructions and high-quality rendering results.
3 Method
3.1 Preliminary
Kerbl et al. [16] proposed representing a scene using 3D Gaussian ellipsoids as primitives and render images using differentiable volume splatting. The shape and orientation of the ellipsoids are controlled by the scale $\mathbf{S}$ and pose $\mathbf{R}$ at their location $\mathbf{p}_{k}$. Together, $\mathbf{R}$, $\mathbf{S}$, and $\mathbf{p}_{k}$ define a metric space, within which a Gaussian distribution is formulated:
$G(\mathbf{x})=\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{p}_{k}))$ (1)
Where the covariance matrix $\mathbf{\Sigma}=\mathbf{RSS^{T}R^{T}}$. To render the 3D Gaussian onto an image, each Gaussian distribution is first transformed into the camera space using the world-to-camera transformation $\mathbf{W}$. It is then mapped into ray space, where the viewing direction is aligned with the coordinate axis, through a local affine transformation $\mathbf{J}$ [39].
$\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$ (2)
Under orthogonal projection, we skip the third row and column of $\mathbf{\Sigma}^{\prime}$, yielding a 2×2 covariance matrix $\mathbf{\Sigma}^{2D}$ to represent the 2D Gaussian $G^{2D}$. Subsequently, 3DGS employs volume rendering to integrate the 2D Gaussian.
$C(\mathbf{p})=\sum_{i=0}^{N-1}G_{i}^{2D}(\mathbf{p})\alpha_{i}c_{i}\prod_{j=0}^{i-1}(1-G_{j}^{2D}(\mathbf{p})\alpha_{j})$ (3)
Here, $\alpha_{i}$ represents the opacity, $c_{i}$ denotes the color of each Gaussian primitive, and $\mathbf{p}$ is the pixel coordinate. Finally, each Gaussian primitive is optimized by minimizing the photometric loss.
GS mesh reconstruction. Vanilla 3DGS [16] can render high-quality images, but it yields suboptimal results for scene geometry reconstruction due to a lack of multi-view consistency in its splatting process. In subsequent geometric reconstruction methods, volumetric approaches like GOF [35] and RadeGS [36] employ ray-splat intersection techniques, achieving state-of-the-art reconstruction quality but limiting the consistency of normal and depth. In contrast, 2DGS [12] defines the 2D Gaussian distribution within a planar disk, which inherently provides consistent normals and depths across multiple views. But the disk is only a first-order approximation of the surface, which often leads to overly smooth reconstruction results in 2DGS, as shown in Fig 2.
3DGS
2DGS
QGS
Figure 2: Demo and mesh reconstruction comparison on Mip-NeRF 360 dataset. 3D Gaussian ellipsoids struggle to fit precisely onto surfaces, resulting in coarse and incomplete meshes. Planes cannot fully capture high-curvature regions, leading to over-smoothed reconstructions. In contrast, quadric surfaces effectively capture geometric edge regions, achieving detailed reconstruction results.
3.2 Quadratic Gaussian splatting
To enhance the geometric fitting capability of the surface representation, we present differentiable Quadratic Gaussian Splatting, as shown in Fig 1. We will first introduce the Quadric Gaussian Model, then discuss the splatting design for quadrics, and finally explain the optimization process.
3.2.1 Quadratic Gaussian Model
Quadratic Model. Given a homogeneous coordinate $\mathbf{x}=[x,y,z,1]^{T}\in\mathbb{R}^{4}$, a quadric surface can be defined as the solution set to the following equation:
$f(x,y,z)$ $=Ax^{2}+2Bxy+2Cxz+2Dx+Ey^{2}$ (4)
$+2Fyz+2Gy+Hz^{2}+2Iz+J$
$=\left[\begin{matrix}x&y&z&1\end{matrix}\right]\left[\begin{matrix}A&B&C&D\\ B&E&F&G\\ C&F&H&I\\ D&G&I&J\end{matrix}\right]\left[\begin{matrix}x\\ y\\ z\\ 1\end{matrix}\right]$
$=\mathbf{x}^{T}\mathbf{Q}\mathbf{x}=0$
Similar to [26], we apply congruent diagonalization to transform the above equation into its canonical form:
$\mathbf{Q}=\mathbf{T}^{-T}\mathbf{D}\mathbf{T}^{-1},\mathrm{with}$ $\mathbf{D}\;\mathrm{diagonal},d_{ii}\in\{0,\pm 1\}$ (5)
$\mathbf{T}=\left[\begin{matrix}\mathbf{u}&\mathbf{v}&\mathbf{w}&\mathbf{c}\\ 0&0&0&1\end{matrix}\right]$
Here, $\mathbf{c}$ denotes the position of the quadric, and $[\mathbf{u},\mathbf{v},\mathbf{w}]$ can be decomposed into $\mathbf{RS}$, where $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}]$ and $\mathbf{S}=\mathrm{diag}(s_{1},s_{2},s_{3})$, which denotes the orientation and scale of the quadric in the object space. The matrix $\mathbf{D}$ defines the surface shape: $\mathbf{D}=\mathrm{diag}(1,1,1,1)$ yields an ellipsoid, while $\mathbf{D}=\mathrm{diag}(1,0,0,0)$ produces a plane.
To compute the Gaussian weight at any surface point, we first define a measure on the surface to establish the Gaussian distribution. To concentrate the Gaussian energy on the surface, we use geodesic length [27] as the metric: the geodesic distance between two surface points is the shortest path along the surface, as illustrated by the red line in Fig 3.
However, not all geodesics have closed-form solutions. When $\mathbf{Q}$ is an ellipsoid or hyperboloid, computing the geodesic length typically requires numerical methods. Therefore, we focus only on the case of a paraboloid:
$f(x,y,z)$ $=\mathbf{x}^{T}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-T}\bar{D}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-1}\mathbf{x}$ (6)
$=\hat{\mathbf{x}}^{T}\left[\begin{matrix}\frac{d_{11}}{s_{1}^{2}}&0&0&0\\ 0&\frac{d_{22}}{s_{2}^{2}}&0&0\\ 0&0&0&-\frac{d_{33}}{2s_{3}}\\ 0&0&-\frac{d_{33}}{2s_{3}}&0\end{matrix}\right]\hat{\mathbf{x}}$
$=\frac{d_{11}}{s_{1}^{2}}\hat{x}^{2}+\frac{d_{22}}{s_{2}^{2}}\hat{y}^{2}-\frac{d_{33}}{s_{3}}\hat{z}=0$
Here and henceforth, we use $\hat{\cdot}$ to denote Gaussian local coordinate. $d_{ii}\in\{0,\pm 1\}$ determines whether the paraboloid is elliptic, hyperbolic, or planar. However, since $d_{ii}$ is discrete, the primitive cannot transition smoothly between elliptic and hyperbolic paraboloids. To resolve this, we introduce a signed scale for a differentiable transition between paraboloid types.
$f(\hat{x},\hat{y},\hat{z})=\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}\hat{x}^{2}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}\hat{y}^{2}-\frac{1}{s_{3}}\hat{z}=0$ (7)
In vanilla 3DGS [16], the scale is obtained through the $\exp$ activation function, i.e., $s(x)=\exp(x)$. To introduce a sign, we add another variable $t$ to control the sign, i.e., $s(x,t)=\tanh(t)\exp(x)$.
Gaussian Distribution on Quadric. We will now describe how to define a Gaussian distribution on a paraboloid. First, we need to define a measure on the quadric. Paraboloid (Equation 7) could be expressed in explicit form:
$\hat{z}(\hat{x},\hat{y})=s_{3}(\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}\hat{x}^{2}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}\hat{y}^{2})$ (8)
We convert it isometrically into cylindrical coordinates, i.e. $\hat{x}=\rho\cos\theta,\hat{y}=\rho\sin\theta$. And we rewrite Equation 8 as:
$\vspace{-5pt}\begin{split}\hat{z}(\theta,\rho)=\;&s_{3}(\frac{\mathrm{sign}(s_{1})\cos^{2}\theta}{s_{1}^{2}}+\frac{\mathrm{sign}(s_{2})\sin^{2}\theta}{s_{2}^{2}})\rho^{2}\\ =\;&a(\theta)\rho^{2}\end{split}\vspace{-5pt}$ (9)
Due to the paraboloid’s symmetry, for any point $\hat{\mathbf{p}}_{0}=(\rho_{0},\theta_{0},\hat{z}(\theta_{0},\rho_{0}))$, the intersection line of the plane $\theta=\theta_{0}$ with the paraboloid: $\hat{z}(\theta_{0},\rho),\rho\in(0,\rho_{0})$, is the geodesic to the origin. Then the geodesic distance is the arc length $l$ of this curve, as shown in Fig 3.
Figure 3: Illustration of a Gaussian on a quadric surface. The blue dot marks the Gaussian centroid, the red dot indicates the ray-splat intersection, and the red line represents the geodesic.
$l(a,\rho_{0})$ $=\int_{0}^{\rho_{0}}\sqrt{1+(2at)^{2}}\mathrm{d}t$ (10)
$=\frac{\ln(\sqrt{u^{2}+1}+u)+u\sqrt{u^{2}+1}}{4a}$
$\mathrm{where}\;u=2a\rho_{0}$
For the derivation of the integral, please refer to the supplementary materials. We then define the mean of the 2D Gaussian distribution on the surface at the origin of the quadric surface, with $(s_{1},s_{2})$ representing the principal axis variances of the Gaussian. Since the contours of the 2D Gaussian distribution form an ellipse:
$\frac{\rho^{2}\cos^{2}\theta}{s_{1}^{2}}+\frac{\rho^{2}\sin^{2}\theta}{s_{2}^{2}}=1$ (11)
Given a point $(\theta_{0},\rho_{0})$ on the ellipse, $\rho_{0}$ represents the standard deviation of the 2D Gaussian distribution in the $\theta_{0}$ direction.
$\sigma(\theta_{0})=\rho_{0}=\frac{s_{1}s_{2}}{\sqrt{(s_{2}\cos\theta_{0})^{2}+(s_{1}\sin\theta_{0})^{2}}}$ (12)
Thus, for any point $\hat{\mathbf{p}}_{0}$ on the surface, we can define the corresponding Gaussian function value as:
$G(\hat{\mathbf{p}}_{0}(\theta_{0},\rho_{0}))=\exp(-\frac{(l(a(\theta_{0}),\rho_{0}))^{2}}{2(\sigma(\theta_{0}))^{2}})$ (13)
Notably, when $|s_{3}|\rightarrow 0$, the paraboloid becomes equivalent to a disk. Furthermore, as $x\rightarrow 0$, we have $\sqrt{1+x}\rightarrow 1$ and $\ln(1+x)\sim x$. Thus, as $|s_{3}|\rightarrow 0$, we get $a\rightarrow 0$ and $l\rightarrow\rho_{0}$ by Equation 10, meaning the geodesic distance becomes equivalent to the Euclidean distance. This indicates that 2DGS can be regarded as a specific degenerate case of QGS, whose more generalized nature allows it to fit high-curvature regions effectively.
3.2.2 Splatting
Although Sigg et al. [26] derived the Ray-Quadric Intersection, QGS integrates Gaussian distribution, necessitating a redefinition of the intersection for multi-view consistency.
Ray-splat Intersection. Let the camera center in the Gaussian local space be denoted as $\hat{\mathbf{o}}\in\mathbb{R}^{3\times 1}$ and the ray direction as $\hat{\mathbf{d}}\in\mathbb{R}^{3\times 1}$. A point on the ray can be defined as $\hat{\mathbf{p}}=\hat{\mathbf{o}}+t\hat{\mathbf{d}}$. By substituting $\hat{\mathbf{p}}$ into the Equation 8, we solve a quadratic equation to obtain two intersection points: the nearer point $\hat{\mathbf{p}}_{n}=(\hat{x}_{n},\hat{y}_{n},\hat{z}_{n})$ and the farther point $\hat{\mathbf{p}}_{f}=(\hat{x}_{f},\hat{y}_{f},\hat{z}_{f})$, with $t_{n}\leq t_{f}$. To ensure multi-view consistency in QGS, we select one of these two points: First, if the geodesic distance of $\hat{\mathbf{p}}_{n}$ is within $3\sigma(\theta_{n})$, we choose $\hat{\mathbf{p}}_{n}$. If not, we check if $\hat{\mathbf{p}}_{f}$ is within $3\sigma(\theta_{f})$. If so, we select $\hat{\mathbf{p}}_{f}$. If neither condition is met, we assume no intersection between the ray $\hat{\mathbf{p}}(t)$ and the primitive. The derivation is provided in the supplementary material.
Normal and Curvature. Similar to 2DGS [12], QGS is a surface-based representation that naturally possesses multi-view consistent geometric properties, making it straightforward to compute surface normals. Given any point $\hat{\mathbf{p}}_{0}=(\hat{x}_{0},\hat{y}_{0},\hat{z}(\hat{x}_{0},\hat{y}_{0}))$ on the surface, we can take the partial derivatives of the Equation 7, yielding:
$\hat{\mathbf{n}}_{0}(\hat{\mathbf{p}}_{0})=(\frac{2\mathrm{sign}(s_{1})}{s_{1}^{2}}\hat{x}_{0},\frac{2\mathrm{sign}(s_{2})}{s_{2}^{2}}\hat{y}_{0},-\frac{1}{s_{3}})$ (14)
As QGS provides a second-order fit, it naturally outputs second-order geometric information like curvature, which describes surface bending. For each QGS primitive, the Gaussian curvature at the ray-splat intersection can be computed analytically, with detailed derivation available in the supplementary material. Let $\lambda_{1}=\mathrm{sign}(s_{1})\cdot s_{3}/s_{1}^{2}$ and $\lambda_{2}=\mathrm{sign}(s_{2})\cdot s_{3}/s_{2}^{2}$. The curvature at point $\hat{\mathbf{p}}_{0}$ is:
$\hat{K}_{0}(\hat{\mathbf{p}}_{0})=\frac{4\lambda_{1}\lambda_{2}}{(1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2})^{2}}$ (15)
We render the normal map $N(\mathbf{p})$ and curvature map $K(\mathbf{p})$ by Equation 16 for a given viewpoint using alpha-blending, as shown in the Fig 4.
$N(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}\textbf{n}_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$ (16)
$K(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}K_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$
normal
curvature
Figure 4: The normal map (left) and curvature map (right) of a scene. In the curvature map, blue indicates higher curvature and curved areas, while red indicates lower curvature and flatter areas.
Per-tile Sorting and Per-pixel Resorting. In Equation 16, the index $i$ is determined by sorting primitives from near to far. Most GS-like methods sort by splat centroid depth rather than ray-splat intersection depth, which, as noted by StopThePop [25], causes popping artifacts. Additionally, we observed that centroid sorting introduces streak artifacts in geometry reconstruction, as shown in Fig 8. To address this, we apply StopThePop’s Per-tile Sorting and Per-pixel Resorting [25] to our QGS. In short, we first compute the ray closest to the quadric center in each tile to estimate intersection depth for rough sorting. During volume rendering, we use a buffer array to locally resort quadrics based on the intersection depth of each pixel. Further details are provided in the supplementary material.
3.3 Optimization
For QGS, which has stronger geometric fitting capability, relying solely on photometric consistency error can introduce geometric noise, such as unsmooth surfaces. Therefore, we adopt the depth distortion loss and normal consistency loss proposed by 2DGS [12] to mitigate this.
Depth Distortion. We employ the depth distortion loss to promote tighter alignment of the quadric surfaces.
$\mathcal{L}_{d}=\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}\omega_{i}\omega_{j}(t_{i}-t_{j})^{2}$ (17)
Here, $\omega_{i}=\bar{\alpha}_{i}T_{i}$ denotes the alpha-blending weight of the i-th Gaussian, and $t_{i}$ represents the depth at the ray-splat intersection. $i,j$ index the Gaussians contributing to the ray. Additionally, following GOF’s recommendation, we freeze the gradient of the depth distortion loss through $\omega_{i}$, optimizing only the depth $t_{i}$.
Normal Consistency. 2DGS [12] introduces the normal consistency loss to ensure that all primitives on a ray are locally aligned with the actual surface.
$\mathcal{L}_{n}=\sum_{i}\omega_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$ (18)
Here, $\mathbf{n}_{i}$ represents the splat normal facing the camera, while $\mathbf{N}$ is computed by differentiating the depth point $\mathbf{p}$ from neighboring pixels.
$\mathbf{N}(u,v)=\frac{\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}}{|\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}|}$ (19)
However, neighboring pixels may not satisfy the local planar assumption, especially in regions with significant depth variation. Using differential normal consistency loss in edge areas can introduce errors, contributing to the over-smoothing observed in 2DGS. PGSR [4] and MVG-splatting [19] have both noted this issue and use image edges to approximate geometric edges, applying additional processing at these edges. We have found that, in most scenes, image edges do not fully correspond to geometric edges, especially in uniformly lit areas. Thus, we utilize the curvature map, which more accurately corresponds to geometric edges and is both efficiently and uniquely generated by QGS, to guide normal supervision.
$\lambda_{K}(K(u,v))=$ $1-\mathrm{sigmoid}(\ln(|K(u,v)|+\varepsilon)$ (20)
$\mathcal{L}_{Kn}(u,v)$ $=\lambda_{K}(K(u,v))\mathcal{L}_{n}(u,v)$
Final Loss. Finally, we input a sparse point cloud and posed images to optimize QGS with the following loss function:
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{Kn}$ (21)
4 Experiment
Figure 5: Qualitative geometric reconstruction comparisons on three datasets (DTU, Tanks and Temples, and Mip-NeRF 360) against 2DGS and GOF, covering indoor, outdoor, object-centric, and forward-facing scenes. QGS shows obvious improvements in both detail and smoothness compared to 2DGS and GOF.
CD (mm)$\downarrow$ 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean Time
NeuS [30] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 $>$12h
VolSDF [32] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 $>$12h
Neuralangelo [18] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 $>$128h
SuGaR[11] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 1h
2DGS [12] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 0.32h
GOF [35] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 1h
Ours 0.46 0.81 0.40 0.42 1.07 0.80 0.79 1.25 0.95 0.67 0.62 1.27 0.40 0.60 0.49 0.73 0.83h
Table 1: Quantitative Chamfer Distance comparison of QGS with GS-like and NeRF-like methods on the DTU dataset. Compared to the planar representation in 2DGS, the surface-based QGS shows a clear improvement in geometric reconstruction quality. Additionally, QGS is an order of magnitude faster than NeRF-like methods.
F1-Score $\uparrow$ NeuS Geo-NeuS Neuralangelo 2DGS GOF Ours
Barn 0.29 0.33 0.70 0.41 0.51 0.43
Caterpillar 0.29 0.26 0.36 0.24 0.41 0.31
Courthouse 0.17 0.12 0.28 0.16 0.28 0.26
Ignatius 0.83 0.72 0.89 0.52 0.68 0.79
Meetingroom 0.24 0.20 0.32 0.17 0.28 0.25
Truck 0.45 0.45 0.48 0.45 0.58 0.60
Mean 0.38 0.35 0.50 0.33 0.46 0.44
Time $>$24h $>$24h $>$127h 0.57h 2h 2h
Table 2: Quantitative F1-Score comparison of QGS with GS-like and NeRF-like methods on the TNT dataset. QGS shows notable gains over 2DGS and is competitive with prior SOTA methods.
We compared our method with several SOTA approaches across multiple datasets, including DTU [14], TNT [17], and Mip-NeRF 360 [2]. We evaluated geometry with F1-score and Chamfer Distance, and appearance with PSNR, SSIM, and LPIPS, followed by analysis and conclusions.
4.1 Implementation Details.
Rasterizer We implemented Quadratic Splatting with custom CUDA kernels on the 3DGS framework, extending the renderer to output depth distortion maps, depth maps, normal maps, and curvature maps. Given the non-convex nature of paraboloids and the lack of an inverse for the geodesic distance function, we used rectangular truncation and approximations to compute image bounding boxes, as detailed in the supplementary materials. To handle the more complex surface intersections, we incorporate Per-tile Sorting, Per-pixel Local Resorting from StopThePop [25], with further details also in the supplementary materials.
Settings. We adopted the adaptive control strategy from 3DGS [16]. Similar to 2DGS [12], the QGS splatting process approximates by projecting the 3D center gradient onto screen space instead of relying directly on the 2D projected center gradient, using a gradient threshold of 0.3 and a percent dense value of 0.001. To ensure a similar number of Gaussian primitives across methods, we also set the gradient threshold for other methods to 2e-4. All experiments were conducted on a single A6000 GPU.
Mesh Extraction. We used a weighted combination of 0.5 median depth (i.e., $t_{\mathrm{median}}=\max{t_{i}|T_{i}>0.5}$) and 0.5 expected depth (i.e., $t_{\mathrm{expected}}=\sum_{i=0}^{N-1}G_{i}\alpha_{i}T_{i}t_{i}$) as the final output depth. Then we fused the depth maps using Truncated Signed Distance Fusion (TSDF) with Open3D [37]. Following the 2DGS setup, we set the voxel size to 0.004 and truncation threshold to 0.002 for fair comparison.
4.2 Comparison
Geometry Evaluation In all experiments, we performed evaluations at half image resolution. In Table 1, we compared QGS with implicit [30, 32, 18] and explicit [12, 35, 11] SOTA reconstruction methods on the DTU dataset using Chamfer Distance as the benchmark. QGS outperformed all explicit methods, and while a gap remains with Neuralangelo [18], our reconstruction time is two orders of magnitude faster. In certain scenes, such as Scan 97 and Scan 105, QGS even outperformed Neuralangelo. In Table 2, we evaluated QGS and other SOTA methods on the TNT dataset using F1-score. In large-scale scenes, QGS consistently outperformed 2DGS [12]. Notably, in the Courthouse and Meetingroom scenes, QGS captured more details due to its higher geometric flexibility, as shown in the Fig 6.
2DGS
QGS
Figure 6: Mesh reconstruction comparison between QGS and 2DGS in large-scale scenes. The first row shows the Courthouse scene, with both using median depth, and the second row shows the Meetingroom scene, using a 0.5 weighted combination of median and expected depth.
As shown in Fig 5, our QGS achieves more complete geometry and stronger detail capture due to the greater flexibility of individual primitives. Note that our mesh extraction is limited by the lower resolution of the voxel grid in TSDF fusion, resulting in slightly lower accuracy compared to GOF [35], which uses marching tetrahedra, leading to longer meshing times and higher VRAM consumption.
QGS’s second-order fit greatly improves on 2DGS [12], capturing finer details and providing second-order geometric information. The curvature map in Fig 4 illustrates how QGS better conforms to object edges with higher curvature, while adhering to flatter surfaces in other areas. It is important to note that the advantage of surface-based representations lies in their ability to provide multi-view consistent depth and normals (and curvature in our method). Thus, both theoretically and practically, it is straightforward to transfer multi-view geometric techniques, as done in PGSR [4] and MVG-Splatting [19], to QGS. We leave this exploration for future work to further enhance our approach.
Rendering Evaluation We compared rendering quality on the Mip-NeRF360 dataset with baseline approaches. As a surface-based method, QGS achieves competitive rendering results in indoor scenes, as shown in Table 3, but performs less effectively in outdoor environments. We attribute this to QGS’s higher geometric fitting capability, which may lead to overfitting in regions with sparse views or low texture. Future work could explore additional regularization constraints, particularly on the curvature of quadric surfaces, to address this issue.
Indoor scenes Outdoor scenes
[…]
NeRF 26.84 0.790 0.370 21.46 0.458 0.515
Deep Blending 26.40 0.844 0.261 21.54 0.524 0.364
i-NGP 29.15 0.880 0.216 22.90 0.566 0.371
Mip-NeRF360 31.72 0.917 0.180 24.47 0.691 0.283
3DGS 30.99 0.926 0.199 24.24 0.705 0.283
SuGar 29.44 0.911 0.216 22.76 0.631 0.349
2DGS 30.39 0.924 0.182 24.33 0.709 0.284
GOF 30.80 0.928 0.167 24.76 0.742 0.225
QGS 30.76 0.926 0.166 24.08 0.699 0.273
Table 3: Quantitative comparison of appearance between QGS, GS-like, and NeRF-like methods on the Mip-NeRF 360 dataset.
4.3 Ablation
Curvature-Guided Normal Consistency Loss. As mentioned in Section 3.2, we use curvature to guide normal consistency supervision. High-curvature regions are treated as high-frequency areas that violate the local planar assumption, so we reduce supervision in these regions. As shown in Table 4 and Fig 7, curvature-guided normal supervision helps recover more scene details, leading to improved geometric quality.
w/o $\lambda_{K}$
w/ $\lambda_{K}$
Figure 7: Comparison of curvature-guided normal consistency versus without curvature guidance. Curvature-guided supervision effectively reduces over-smoothing and recovers more details.
F1-Score $\uparrow$ Truck Ignatius
[…]
w/o $\lambda_{K}$ 0.63 0.51 0.57 0.74 0.75 0.75
w/ $\lambda_{K}$ 0.65 0.55 0.60 0.78 0.80 0.79
Table 4: Quantitative F1-Score comparison on the TNT dataset with and without curvature-guided normal supervision.
Per-tile Sorting. 2DGS suggests the actual surface lies at the median intersection point where opacity reaches 0.5. However, most GS-like methods sort Gaussians by centroid depth, leading to streak-like inconsistencies, as shown in Fig 8. To address this, we implemented per-tile sorting (TS) and per-pixel resorting (PR) based on the methods proposed in StopThePop [25], further refining them to remove inconsistencies and enhance reconstruction quality, as shown in Table 5.
w/o TS and PR
w/ TS
w/ TS and PR
Figure 8: Normal comparison with and without per-tile sorting. Centroid depth sorting causes noticeable streak-like inconsistencies, also seen in 2DGS.
CD (mm)$\downarrow$ 24 37 97 122
w/o TS and PR 0.48 0.85 1.04 0.55
w/ TS 0.46 0.80 0.96 0.52
w/ TS + PR 0.46 0.81 0.95 0.49
Table 5: Quantitative Chamfer Distance comparison on the DTU dataset with and without per-tile sorting.
5 Conclusion
In this work, we introduce Quadratic Gaussian Splatting, a variant of GS-like methods, designed to reconstruct accurate scene geometry and recover finer details. QGS is the first to introduce quadric surfaces to Gaussian Splatting, defining Gaussian distributions in non-Euclidean space to improve fitting and capture second-order curvature. We achieve SOTA geometric reconstruction and competitive rendering results on various indoor and outdoor datasets. Additional results are in the supplementary materials.
Discussion. Gaussian Splatting is a rapidly evolving field, with numerous works exploring its use for surface reconstruction, all developed independently from ours, such as Rade-GS [36], PGSR [4], Trim-GS [7], and MVG-Splatting [19]. Our approach differs in focus, introducing a novel representation that provides curvature information. Our contributions are orthogonal to these approaches and, in theory, could enhance their methods as well.
Limitation. Since the bounding box of QGS primitives on the image is approximated, QGS processes more tiles than typical GS methods, particularly for larger quadric surfaces, resulting in a speed disadvantage. As a surface-based representation, QGS is still limited in capturing foggy scenes or transparent objects.
Supplementary Material
In this supplementary material, we provide detailed explanations of the following: (1) Calculation of the projected bounding box of primitives. (2) Details of per-tile and per-pixel sorting. (3) Solving the ray-primitive intersection equation and numerical handling. (4) Derivation of the geodesic distance formula integration. (5) Derivation of Gaussian curvature. (6) We present more qualitative results. Additionally, we include a video (QGS.mp4) that briefly explains the main workflow of QGS and showcases comparative results.
Details of Bounding Box Calculation.
In 2DGS/3DGS [12, 16], Gaussian primitives are enclosed convex representations, making it straightforward and convenient to compute their bounding box on the image, as shown in the first column of Fig 9.
Figure 9: Comparison diagram of the bounding boxes between 2DGS and QGS.
However, QGS includes concave cases, and geodesic distance complicates analytical determination of the Gaussian primitive’s rendered portion during preprocessing. Specifically, the geodesic function Equation 10 in the main text lacks an elementary inverse, so computing the equipotential $\rho(\theta_{0})=\{l^{-1}(l_{0}):l_{0}=\sigma(\theta_{0})\}$, requires numerical or approximate solutions. Even with an analytical solution for ${\rho(\theta_{0})=l^{-1}(\sigma(\theta_{0}))}$, the Gaussian distribution in non-Euclidean space means equipotential lines may not fully enclose the primitive, requiring extra boundary calculations. To simplify, we use the quadric’s circumscribing rectangular box as the 3D bounding box. Its 8 vertices are projected onto the image plane, and their 2D AABB is used as the 2D bounding box, as shown in Fig 9, columns two and three. Specifically, we first scale the geodesic function to a quadratic polynomial function:
$l(a,\rho)\approx\hat{l}(a,\rho)=\frac{4|a|\rho^{2}+6\rho}{7}$ (22)
Given the direction $\theta_{0}$, we obtain the variance of the Gaussian distribution $\sigma_{0}(\theta_{0})$. The root of the equation $\hat{l}(a,\rho)-\sigma_{0}=0$ in this direction can be calculated as:
$\rho_{0,1}=\frac{-6\pm\sqrt{36+112|a|\sigma_{0}}}{8|a|}$ (23)
The comparison between the approximate function and the geodesic distance function is shown in Fig 10, demonstrating that the two functions nearly overlap.
$a=0.01$
$a=0.1$
$a=0.5$
Figure 10: Comparison diagram of the approximate function versus the geodesic function. The red line represents the approximate function, while the black line represents the geodesic function.
Since the quadratic curve passes through the origin, we take its positive root, representing the horizontal distance from the primitive’s vertex at a geodesic distance of $\sigma_{0}$. We compute this for the major and minor axes, i.e. $\theta=0$ and $\theta=\pi/2$, yielding $\rho_{l}$ and $\rho_{s}$, with circumscribing box radii $(\rho_{l},\rho_{s},a\max(\rho_{l}^{2},\rho_{s}^{2}))$. We then project the eight vertices of the bounding box onto the image and take the minimum 2D bounding box of the resulting polygon. Due to approximations, gaps may appear between the bounding box and the primitive as shown in the third column of Fig 9, affecting rendering speed but not quality.
Details of Per-tile Sorting and Per-pixel Resorting
2DGS [12] suggests the surface lies at the median intersection point where opacity reaches 0.5. However, for surface-based representations like 2DGS and QGS, the Gaussian distribution is concentrated on the surface, making the alpha-blending order more sensitive. Most reconstruction methods [35, 36, 19, 4], including 2DGS, sort by Gaussian centroid depth, causing geometric inconsistencies, as shown in Fig 8 of the main text.
To address this, we introduce StopThePop’s per-tile sorting and per-pixel resorting [25] for more precise ordering. Specifically, for each $16\times 16$ pixel tile, we compute the intersection depth using the ray from the pixel closest to the projected vertex of the quadric surface and apply this depth to all 256 pixels for tile-based global sorting. As shown in Fig 8 in the main text, this method effectively removes streak-like inconsistencies but introduces minor blocky artifacts due to approximating with a single ray per tile. Then we adopt StopThePop’s per-pixel local resorting to reorder the Gaussians along each ray, to eliminate blocky inconsistencies. Specifically, after calculating each Gaussian’s depth, normal, and other properties, we do not use them for alpha-blending immediately. Instead, we store them in an 8-length buffer array, and once the buffer is full, we select the closest Gaussian for alpha-blending. In original 3DGS, forward and backward computations use different traversal orders, causing the buffering mechanism to produce inconsistent rendering. Thus, following StopThePop’s strategy, we perform gradient computation in a near-to-far order, which requires modifying the original gradient formulas. For volumetric rendering maps, we can uniformly express them as:
$\hat{X}=\sum_{i=0}^{N-1}\bar{\alpha}_{i}T_{i}X_{i},\;\;\mathrm{where}\;\bar{\alpha}_{i}=\alpha_{i}g_{i},T_{i}=\prod_{j=0}^{i-1}(1-\bar{\alpha}_{j})$
Where $X_{i}$ represents properties such as the color, depth, normal, or curvature of the primitives. The gradient computation in a back-to-front order is given as follows:
$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}$ $=X_{i}T_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{1-\bar{\alpha}_{i}}$
$=(X_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
We rewrite it as front-to-back form:
$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}=(X_{i}-\frac{\hat{X}-\sum_{j=0}^{i}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
Ensuring that the backward computation follows the same order as the forward computation. Fortunately, the gradient computation order does not affect the distortion loss, meaning no additional derivation is needed for the distortion loss.
Figure 11: Illustration of approximate intersections. The blue thick line represents the depth of the approximate intersection.
Details of Ray-splat Intersection
Given the camera center $\hat{\mathbf{o}}=[\hat{o}^{1},\hat{o}^{2},\hat{o}^{3}]^{T}$ and ray direction $\hat{\mathbf{r}}=[\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}]^{T}$ in the Gaussian coordinate system, the ray can be expressed as:
$\hat{\textbf{x}}=\left[\begin{matrix}\hat{o}^{1}+t\hat{r}^{1}\\ \hat{o}^{2}+t\hat{r}^{2}\\ \hat{o}^{3}+t\hat{r}^{3}\end{matrix}\right]$ (24)
Here, $t$ denotes the depth. By solving the ray equation 24 together with the quadric surface equation 7 in the main text, we find two intersection points:
$t^{2}[\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}(\hat{r}^{1})^{2}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}(\hat{r}^{2})^{2}]+$ (25)
$t[\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}2\hat{o}^{1}\hat{r}^{1}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}2\hat{o}^{2}\hat{r}^{2}-\frac{1}{s_{3}}\hat{r}^{3}]+$
$[\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}(\hat{o}^{1})^{2}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}(\hat{o}^{2})^{2}-\frac{1}{s_{3}}\hat{o}^{3}]=0$
$At^{2}+Bt+C=0$
$t_{n}=\frac{-B-\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A},$
$t_{f}=\frac{-B+\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A}$
However, Equation 25 encounters numerical issues when $|A|$ is very small, particularly during backpropagation, as a small denominator can cause floating-point overflow, resulting in invalid values such as NaN or Inf. Upon examination, we observe that when $A$ is especially small, it corresponds to cases where the quadric surface is nearly flat or when the ray is almost perpendicular to the $s_{1}\times s_{2}$ plane. In such cases, we can ignore the quadratic term, reducing Equation 25 to $Bt+C=0$, or $t=-\frac{C}{B}$. Geometrically, this solution represents the depth of the intersection between the ray and the tangent plane determined by the point where the perpendicular line from the camera center meets the quadric surface, as indicated by the thick blue line $n$ in Fig 11.
Let $\mathbf{M}=s_{1}\times s_{2}$ be the horizontal plane in the Gaussian local coordinate system. Let $\mathbf{p}_{o}=[\hat{o}_{1},\hat{o}_{2},s_{3}(\frac{\hat{o}_{1}^{2}}{s_{1}^{2}}+\frac{\hat{o}_{2}^{2}}{s_{2}^{2}})]^{T}$ represent the intersection of the perpendicular line from the camera center to $\mathbf{M}$ and the quadric surface. The normal of the tangent plane $\mathbf{N}$ at the intersection point can be expressed as:
$\mathbf{\hat{n}}(\mathbf{p}_{o})=[\frac{2\mathrm{sign}(s_{1})\hat{o}_{1}}{s_{1}^{2}},\frac{2\mathrm{sign}(s_{2})\hat{o}_{2}}{s_{2}^{2}},-\frac{1}{s_{3}}]^{T}$
The projection of $(\mathbf{p}_{o}-\hat{\mathbf{o}})$ onto the normal direction $\mathbf{n}$ is:
$m=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{||\hat{\mathbf{n}}(\mathbf{p}_{o})||}$
As indicated by the thick red line segment in the Fig 11. On the other hand, the cosine of the angle between the normal vector and the viewing direction can be expressed as:
$\mathrm{cos}\theta=\frac{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}{||\mathbf{\hat{r}}||\cdot||\mathbf{\hat{n}}(\mathbf{p}_{o})||}$
Noting that $||\hat{\mathbf{r}}||=1$, the length of the line from the camera center along the viewing direction to the intersection with the tangent plane is given by:
$n$ $=\frac{m}{\cos\theta}=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}$
$=-\frac{C}{B}$
It is evident that when the surface is nearly flat or the viewing direction is almost perpendicular to the horizontal plane, the approximate solution nearly coincides with the exact solution. Therefore, for cases where $|A|<1\mathrm{e}-6$, we approximate the solution using the method described above.
Derivation of Geodesic Arc Length
For the geodesic arc length formula:
[…]
Let $u=2at,x=\arctan(u)$, then Equation 26 becomes:
[…]
$=\frac{1}{2a}\int\sqrt{1+\tan^{2}(x)}d\tan(x)$
$=\frac{1}{2a}\int\sec(x)d\tan(x)$
[…]
Equation 27 can be solved using integration by parts:
[…]
$=(u\sqrt{1+u^{2}}+\ln(u+\sqrt{1+u^{2}}))/2$
$\Rightarrow l(a,\rho_{0})$ $=\frac{u\sqrt{1+u^{2}}+\ln(u+\sqrt{1+u^{2}})}{4a}$
Details of Curvature
Here, we compute the Gaussian curvature analytically using a standard differential geometry approach [27]. Given the intersection point $\hat{\mathbf{p}}_{0}=[\hat{x}_{0},\hat{y}_{0},\hat{z}_{0}]^{T}$, we simplify Equation 7 as $\hat{z}=\lambda_{1}\hat{x}^{2}+\lambda_{2}\hat{y}^{2}$. The partial derivatives at $\hat{\mathbf{p}}_{0}$ are:
$x_{u}$ $=(1,0,2\lambda_{1}\hat{x}_{0})$
$x_{v}$ $=(0,1,2\lambda_{2}\hat{y}_{0})$
The first fundamental form is:
$E=<x_{u},x_{u}>$ $=1+4\lambda_{1}^{2}\hat{x}_{0}^{2}$
$F=<x_{u},x_{v}>$ $=4\lambda_{1}\lambda_{2}\hat{x}_{0}\hat{y}_{0}$
$G=<x_{v},x_{v}>$ $=1+4\lambda_{2}^{2}\hat{y}_{0}^{2}$
The second fundamental form is:
$n=\frac{x_{u}\times x_{v}}{||x_{u}\times x_{v}||}$ $=\frac{(-2\lambda_{1}\hat{x}_{0},-2\lambda_{2}\hat{y}_{0},1)}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
$x_{uu}$ $=(0,0,2\lambda_{1})$
$x_{uv}$ $=(0,0,0)$
$x_{vv}$ $=(0,0,2\lambda_{2})$
$L=<n,x_{uu}>$ $=\frac{2\lambda_{1}}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
$M$ $=<n,x_{uv}>=0$
$N=<n,x_{vv}>$ $=\frac{2\lambda_{2}}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
Finally, the Gaussian curvature can be computed as:
$K=\frac{LN-M^{2}}{EG-F^{2}}=\frac{\frac{4\lambda_{1}\lambda_{2}}{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}$
Addtional Results. In this section, we present additional qualitative results of QGS in both indoor and outdoor scenarios. Fig 12 compares QGS and 2DGS on the DTU dataset, where QGS demonstrates sharper contours. QGS also achieves accurate reconstructions in indoor and outdoor scenes, as shown in Fig 13 with results from the TNT and Mip-NeRF 360 datasets. Rendering results for all three datasets are shown in Fig 14. Finally, we qualitatively tested our method on an urban scene captured from aerial view. As shown in Fig 15, QGS captures more building details than 2DGS [12] in aerial views. Additionally, most methods using the default 3DGS configuration failed to reconstruct the Courthouse scene and the urban scene, both containing nearly 1,000 images. For these two scenes, we simply doubled the training parameters, setting the total iterations to 60,000, densifying every 800 iterations from the 4,000th to the 48,000th iteration, and resetting Gaussian opacity every 6,000 iterations. For other scenes, we used the default 3DGS configuration.
Figure 12: Qualitative geometry comparison for the DTU dataset [14].
Figure 13: Qualitative geometry results for the Tanks and Temples dataset [17] and Mip-NeRF 360 dataset [2].
Figure 14: Qualitative appearance results for the DTU, TNT and Mip-NeRF 360 dataset.
Figure 15: Qualitative geometry result for the urban scene captured from aerial view, involving over 1,000 images. In the subfigure, the left shows the results of 2DGS, while the right shows the results of QGS.

### v3

@@
-Quadratic Gaussian Splatting for Efficient and Detailed Surface Reconstruction
+Quadratic Gaussian Splatting: High Quality Surface Reconstruction with Second-order Geometric Primitives
 Abstract
-Recently, 3D Gaussian Splatting (3DGS) has attracted attention for its superior rendering quality and speed over Neural Radiance Fields (NeRF).
-To address 3DGS’s limitations in surface representation, 2D Gaussian Splatting (2DGS) introduced disks as scene primitives to model and reconstruct geometries from multi-view images, offering view-consistent geometry.
-However, the disk’s first-order linear approximation often leads to over-smoothed results.
-We propose Quadratic Gaussian Splatting (QGS), a novel method that replaces disks with quadric surfaces, enhancing geometric fitting.
-QGS defines Gaussian distributions in non-Euclidean space, allowing primitives to capture more complex textures.
-As a second-order surface approximation, QGS also renders spatial curvature to guide the normal consistency term, to effectively reduce over-smoothing.
-Moreover, QGS is a generalized version of 2DGS that achieves more accurate and detailed reconstructions, as verified by experiments on DTU and TNT, demonstrating its effectiveness in surpassing most state-of-the-art methods in geometry reconstruction.
-Our code will be released as open source.
-The project is available at https://quadraticgs.github.io/QGS
-Figure 1: Quadratic Gaussian Splatting, a novel quadric-based approach for recovering high-precision real-world geometry from multi-view RGB images.
-In the Courthouse scene, our method achieves multi-view consistent normals, depth, and curvature, producing high-quality meshes and realistic images.
-Compared to 2DGS, QGS more accurately models the building’s contours.
+We propose Quadratic Gaussian Splatting (QGS), a novel representation that replaces static primitives with deformable quadric surfaces (e.g., ellipse, paraboloids) to capture intricate geometry.
+Unlike prior works that rely on Euclidean distance for primitive density modeling—a metric misaligned with surface geometry under deformation—QGS introduces geodesic distance-based density distributions.
+This innovation ensures that density weights adapt intrinsically to the primitive curvature, preserving consistency during shape changes (e.g., from planar disks to curved paraboloids).
+By solving geodesic distances in closed form on quadric surfaces, QGS enables surface-aware splatting, where a single primitive can represent complex curvature that previously required dozens of planar surfels, potentially reducing memory usage while maintaining efficient rendering via fast ray-quadric intersection.
+Experiments on DTU, Tanks and Temples, and MipNeRF360 datasets demonstrate state-of-the-art surface reconstruction, with QGS reducing geometric error (chamfer distance) by 33% over 2DGS and 27% over GOF on the DTU dataset.
+Crucially, QGS retains competitive appearance quality, bridging the gap between geometric precision and visual fidelity for applications like robotics and immersive reality.
+GT
+2DGS
+PGSR
+QGS (8K)
+Figure 1: Our QGS model employs quadric surfaces to recover high-precision geometry from multi-view RGB images, reducing point count while enhancing accuracy.
+To isolate geometry from texture, we evaluate QGS on a weakly textured, geometry-rich subset of the Google Scanned Object dataset [9].
+The results suggest that QGS, under different gradient threshold settings (QGS-x), consistently outperforms 2DGS [16] and achieves comparable quality to PGSR [5] with fewer points.
 1 Introduction
-Surface reconstruction and realistic novel view synthesis (NVS) are critical tasks in computer graphics and vision.
-Their goal is to recover dense geometric structures from images captured from different viewpoints and render photorealistic images.
-Recently, 3D Gaussian Splatting (3DGS) has surpassed Neural Radiance Fileds (NeRF) based methods [2, 23, 24] in both rendering quality and speed by combining traditional splatting techniques with end-to-end optimization.
-Subsequently, GS-based methods have seen rapid advancements in dynamic reconstruction [13, 8], editing [6, 29], and large-scale scene reconstruction [20, 21].
-However, vanilla splatting ignores the Gaussian contribution along the z-axis and uses approximation [39, 16].
-As a result, it can render high-quality, multi-view consistent textures even at offsets from the scene surface, making it challenging to capture accurate geometry.
-In subsequent work [35, 19, 12, 4, 36], ray-splat intersections are computed in 3D to retrieve Gaussian weights, a key factor for accurate scene geometry recovery.
-In methods with Gaussian ellipsoids as primitives [35, 36], ray-splat intersections depend on the view direction, limiting normal consistency across views.
-In contrast, methods using Gaussian disks as primitives [12, 19] can ensure multi-view consistent geometry due to their surface-based nature, facilitating the incorporation of multi-view geometric consistency [4, 19].
-However, disks, as a first-order fit of geometry, often result in overly smooth reconstructions, as shown in the Fig 2.
-Earlier works [26, 40, 31] demonstrated that surfel elements can effectively represent complex scene geometry, such as atomic models.
-Building on this, 2DGS [12] combined Gaussian distributions with disk elements to achieve state-of-the-art (SOTA) reconstruction, inspiring our approach.
-In this paper, we propose a novel surface-based rendering method, which we call Quadratic Gaussian Splatting (QGS), for 3D scene reconstruction and novel view synthesis.
-This method enhances the geometric fitting of scene primitives through higher-order representation, enabling the extraction of accurate surface mesh from the quadric model.
-Unlike previous splatting techniques operating in Euclidean space, such as 3DGS or 2DGS, our QGS defines the Gaussian distribution on a quadratic paraboloid, which can continuously transition between convex and concave forms, offering greater geometric fitting flexibility.
-The key technique in using paraboloids for splatting lies in establishing Gaussian distributions based on geodesic distances in non-Euclidean space, which allows the distribution’s energy to be concentrated on the surface, to effectively capture complex geometric textures.
-Compared to the first-order approximation in 2DGS, QGS provides a second-order fit to the scene geometry.
-As a result, QGS not only produces multi-view consistent normals and depth but also provides curvature information, which represents surface bending and helps guide single-view normal consistency supervision.
-Furthermore, we observe that the volumetric rendering order used by most GS methods causes popping artifacts, which impact both novel view synthesis and geometric reconstruction.
-To address this, we adopted the sorting criteria from StopThePop [25], restructuring it to accommodate quadric surface structures, thereby improving the centroid sorting used in 2DGS to better handle complex intersections of quadratic surfaces.
-Finally, QGS is able to extract high-quality and detail-rich mesh models, as shown in the Fig 1.
-In summary, the main contributions of our work are as follows:
-  • We propose QGS, a novel and efficient differentiable representation that uses quadric surfaces as scene primitives, offering stronger geometric fitting capabilities.
-  • We propose the first Gaussian Splatting work to introduce geodesic distance to establish Gaussian distributions on surfaces, enabling primitives to fit more complex textures.
-  • With stricter depth sorting in the higher-degree-of-freedom QGS, our method achieves SOTA geometric reconstruction and enhanced rendering quality.
-2 Related work
-2.1 Novel View Synthesis
-Novel view synthesis has seen rapid development in recent years, especially since the introduction of NeRF [23].
-NeRF encodes scene geometry and view-dependent appearance using a multi-layer perceptron (MLP) and optimizes it end-to-end through volumetric rendering to produce realistic images.
-Subsequent methods introduced various improvements.
-Mip-NeRF [1], Mip-NeRF 360 [2], and Zip-NeRF [3] addressed NeRF’s aliasing issues by introducing new sampling strategies.
-I-NGP [24] and DVGO [28] significantly accelerated NeRF’s training and rendering speeds by encoding scenes into feature grids instead of MLPs.
-Recently, 3DGS [16] has gained attention for its more realistic rendering capabilities and real-time rendering speed, surpassing NeRF in many cases [34, 25, 22, 38].
-Mip-Splatting [34] introduced a 3D smoothing filter to eliminate high-frequency artifacts during rendering.
-StopThePop [25] proposed per-tile and per-pixel resorting to resolve multi-view geometric inconsistencies.
-Scaffold-GS [22] effectively reduced redundant Gaussians and improved rendering quality by constructing anchor points to distribute local 3D Gaussians.
-2.2 Neural Surface Reconstruction
-Compared to traditional multi-view geometric reconstruction, neural volumetric rendering methods often produce smoother and more complete results.
-NeuS [30] and VolSDF [32] introduced signed distance fields (SDF) to describe scene geometry based on NeRF [23].
-Geo-NeuS [9] and NeuralWarp [10] improved reconstruction by incorporating multi-view consistency.
-Neuralangelo [18] reduced MLP reliance by using feature grids and numerical gradients to improve geometric reconstruction quality.
-However, neural rendering-based surface reconstruction generally requires several hours to converge using only images.
-Additionally, due to the structured implicit fields, accessing and editing scene geometry is not straightforward.
-2.3 Gaussian Splatting Surface Reconstruction
-With the rapid development of 3DGS across various fields, Gaussian splatting methods have also made significant progress in surface reconstruction.
-SuGaR [11] introduced regularization terms to encourage Gaussians to fit the scene surface by building volumetric density fields from Gaussian ellipsoids and using Poisson reconstruction [15] to obtain a mesh.
-GSDF [33] and NeuSG [5] combined 3D Gaussian ellipsoids with SDF, achieving high-quality geometric reconstructions and rendering results.
-Rade-GS [36], GOF [35], and PGSR [4] computed ray-Gaussian ellipsoid intersections to obtain unbiased depth and used the normal consistency supervision from 2DGS [12] to achieve state-of-the-art reconstruction results.
-2DGS flattened Gaussian ellipsoids into Gaussian disks to better align Gaussian primitives with surfaces and introduced two regularization losses to provide additional geometric constraints.
-MVG-splatting [19] extended 2DGS by introducing multi-view consistency constraints, enabling more complex surface reconstructions.
-In this paper, we propose Quadratic Gaussian Splatting (QGS), a new surface representation based on quadric surfaces designed to improve the local geometric fitting capability of primitives.
-We further establish Gaussian distributions on quadric, enabling end-to-end optimization.
-Lastly, we address the volume rendering order issue in 2DGS by incorporating the sorting method from StopThePop [25], achieving precise geometric reconstructions and high-quality rendering results.
-3 Method
-3.1 Preliminary
-Kerbl et al.
-[16] proposed representing a scene using 3D Gaussian ellipsoids as primitives and render images using differentiable volume splatting.
-The shape and orientation of the ellipsoids are controlled by the scale $\mathbf{S}$ and pose $\mathbf{R}$ at their location $\mathbf{p}_{k}$.
-Together, $\mathbf{R}$, $\mathbf{S}$, and $\mathbf{p}_{k}$ define a metric space, within which a Gaussian distribution is formulated:
-$G(\mathbf{x})=\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{p}_{k}))$ (1)
-Where the covariance matrix $\mathbf{\Sigma}=\mathbf{RSS^{T}R^{T}}$.
-To render the 3D Gaussian onto an image, each Gaussian distribution is first transformed into the camera space using the world-to-camera transformation $\mathbf{W}$.
-It is then mapped into ray space, where the viewing direction is aligned with the coordinate axis, through a local affine transformation $\mathbf{J}$ [39].
-$\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$ (2)
-Under orthogonal projection, we skip the third row and column of $\mathbf{\Sigma}^{\prime}$, yielding a 2×2 covariance matrix $\mathbf{\Sigma}^{2D}$ to represent the 2D Gaussian $G^{2D}$.
-Subsequently, 3DGS employs volume rendering to integrate the 2D Gaussian.
-$C(\mathbf{p})=\sum_{i=0}^{N-1}G_{i}^{2D}(\mathbf{p})\alpha_{i}c_{i}\prod_{j=0}^{i-1}(1-G_{j}^{2D}(\mathbf{p})\alpha_{j})$ (3)
-Here, $\alpha_{i}$ represents the opacity, $c_{i}$ denotes the color of each Gaussian primitive, and $\mathbf{p}$ is the pixel coordinate.
-Finally, each Gaussian primitive is optimized by minimizing the photometric loss.
-GS mesh reconstruction.
-Vanilla 3DGS [16] can render high-quality images, but it yields suboptimal results for scene geometry reconstruction due to a lack of multi-view consistency in its splatting process.
-In subsequent geometric reconstruction methods, volumetric approaches like GOF [35] and RadeGS [36] employ ray-splat intersection techniques, achieving state-of-the-art reconstruction quality but limiting the consistency of normal and depth.
-In contrast, 2DGS [12] defines the 2D Gaussian distribution within a planar disk, which inherently provides consistent normals and depths across multiple views.
-But the disk is only a first-order approximation of the surface, which often leads to overly smooth reconstruction results in 2DGS, as shown in Fig 2.
+The pursuit of photorealistic view synthesis and geometry reconstruction is a central focus in academia and industry, with applications spanning virtual reality, film-making, and autonomous driving.
+Recent point-based methods, such as 3D Gaussians [21, 50], have proven highly effective, delivering view synthesis quality comparable to Neural Radiance Fields (NeRF)-based approaches while preserving superior geometry [16, 46] and offering significantly faster performance.
+Since then, these methods have been extended to dynamic reconstruction [17, 10, 27], scene editing [7, 41, 40, 54], and large-scale scene reconstruction [26, 28, 44, 29].
+However, due to their approximation-based nature, 3D Gaussian methods often face challenges like high memory consumption, blurriness, and geometric inaccuracies.
+To address these limitations, recent efforts have focused on “primitive-level modifications.” Methods like GES [14] generalize density distributions to reduce point counts, while others employ sharp kernels (e.g., clipped Gaussian [23], radial [18], or convex kernels [15]) to capture high-frequency appearance details.
+Assigning texture maps to planar primitives has also enabled more efficient scene representations [34, 39, 36, 4, 43].
+Despite these advances, such modifications mainly target high-quality appearance reconstruction, leaving detailed geometry reconstruction largely unaddressed.
+Several approaches adopt alternative kernel shapes, such as flat ellipsoids or planar disks (e.g., surfels) [46, 16, 8] to improve surface alignment.
+Yet, as a first-order geometric approximation, planar disks lack sufficient expressiveness; For example, highly curved surfaces requires dense primitive sampling, creating memory bottlenecks and rendering latency.
+In this paper, we extend the representations of ellipsoid [21] and elliptical disks to a more general quadratic framework.
+Specifically, we define primitive boundaries using an implicit function $f(x,y,z)=0$ where $f(x,y,z)<0$ for points inside the primitive and $f(x,y,z)>0$ otherwise.
+By adopting a quadratic form $f(x,y,z)=(x,y,z,1)^{\top}\mathbf{Q}(x,y,z,1)=0$, where $\mathbf{Q}\in\mathbb{R}^{4\times 4}$ is a conic matrix, we allow the representation of various shapes, including cylinder , ellipsoids, paraboloids, and hyperboloids.
+As a result, it generalizes 3D Gaussian primitives [21] and includes 2D Gaussian splatting [16] as a special case, allowing higher-order surface adaptation while incorporating kernel-level innovations such as shape kernels [18], texture billboards [43, 34, 39], and weight distributions [14].
+In particular, there is an efficient ray-quadratic intersection algorithm [35] that makes this representation practical and versatile in the context of splatting.
+Since the primitive is essentially a second-order primitive, we term our method Quadratic Gaussian Splatting (QGS).
+However, for deformable quadratic primitives, defining a continuous density distribution is a key challenge.
+Unlike methods using rigid primitives (e.g., triangles, spheres) with Euclidean-based density, our deformable quadratic surfaces require a geometry-aware density metric.
+For instance, as Fig 4 shows, when a primitive deforms from a disk to a paraboloid, Euclidean distance would measure straight-line distance, which ignores the curved surface.
+While Geodesic distance follows the surface, yields more accurate density modeling, it is generally difficult to compute.
 3DGS
@@
 QGS
-Figure 2: Demo and mesh reconstruction comparison on Mip-NeRF 360 dataset.
-3D Gaussian ellipsoids struggle to fit precisely onto surfaces, resulting in coarse and incomplete meshes.
-Planes cannot fully capture high-curvature regions, leading to over-smoothed reconstructions.
+Figure 2: Demo and mesh reconstruction comparison on the Mip-NeRF 360 dataset.
+ellipsoids struggle to align accurately with surfaces, producing coarse and incomplete meshes.
+Planes fail to capture high-curvature regions, resulting in oversmoothed reconstructions.
 In contrast, quadric surfaces effectively capture geometric edge regions, achieving detailed reconstruction results.
+Luckily, we target at surface modeling so we constrain our quadratic representation to paraboloids—including planar ellipses as degenerate cases—where efficient closed-form geodesic solvers exist.
+This formulation maintains point-based rendering efficiency while enabling adaptive curvature modeling, where a single paraboloid approximates complex geometry that otherwise requires multiple planar surfels, achieving higher fidelity with similar primitive counts.
+With such surfel-based representation and inspired by 2DGS [16], we adopt a ray-splat intersection [35] for splatting, maintaining multi-view consistency and perspective correctness.
+The surfel nature also allow QGS to leverage advanced splatting techniques that address perspective distortion [16] and popping effects [33], or advanced surface regularization [16, 2, 11] that improves overall geometry accuracy, albeit with added computational cost.
+To validate the effectiveness, we evaluate QGS on three benchmark datasets—MipNeRF360 [2], DTU [19], and Tanks and Temples [22]—and demonstrate state-of-the-art surface reconstruction accuracy.
+Our ablation studies reveal that deformable primitives improve geometric fidelity by 33% over 2DGS [16], establishing QGS as a robust alternative for applications demanding higher geometric precision.
+We summarize our key contributions as follows:
+  • A quadric surfel representation that generalizes 2D Gaussians to higher-order surfaces (e.g., paraboloids), enabling adaptive curvature modeling and capturing high frequency details.
+  • Geodesic-aware density distributions, the first method to unify deformable primitives with surface-aligned density modeling via closed-form geodesic solvers.
+  • State-of-the-art reconstruction: QGS reduces geometric error by 33% over 2DGS and 27% over GOF on DTU [19] while maintaining competitive training and rendering speeds.
+2 Related work
+2.1 Novel View Synthesis
+Novel view synthesis has advanced rapidly, particularly since NeRF [31].
+NeRF represents scene geometry and view-dependent appearance using a multi-layer perceptron (MLP) and optimizes it end-to-end through volumetric rendering to generate realistic images.
+Subsequent methods have introduced various enhancements.
+Mip-NeRF [1], Mip-NeRF 360 [2], and Zip-NeRF [3] addressed aliasing issues by refining sampling strategies, while I-NGP [32] and DVGO [38] significantly improved training and rendering speeds by replacing MLPs with feature grids.
+More recently, 3DGS [21] has gained attention for its superior rendering quality and real-time performance, often surpassing NeRF [49, 33, 30, 54].
+Mip-Splatting [49] introduced a smoothing filter to reduce high-frequency artifacts, while StopThePop [33] improved multi-view consistency through per-tile and per-pixel sorting.
+Scaffold-GS [30] optimized Gaussian distributions by reducing redundancy and improving rendering quality through anchor-based 3D Gaussian placement.
+2.2 Neural Surface Reconstruction
+Neural volumetric rendering methods often produce smoother and more complete reconstructions than traditional approaches.
+NeuS [42] and VolSDF [45] introduced signed distance fields (SDF) to model scene geometry within the NeRF framework [31].
+Geo-NeuS [11] and NeuralWarp [12] further improved reconstruction by enforcing multi-view consistency.
+Neuralangelo [24] reduced reliance on MLPs by incorporating feature grids and numerical gradients, enhancing geometric fidelity.
+However, neural rendering-based surface reconstruction often takes hours to converge using only image supervision.
+Additionally, the structured implicit representations make direct access and editing of scene geometry challenging.
+2.3 Gaussian Splatting Surface Reconstruction
+With the rapid advancement of 3DGS across various domains, Gaussian splatting methods have also significantly improved surface reconstruction.
+SuGaR [13] introduced regularization terms to encourage Gaussians to align with scene surfaces, constructing volumetric density fields from Gaussian ellipsoids and applying Poisson reconstruction [20] to extract meshes.
+GSDF [48] and NeuSG [6] combined 3D Gaussian ellipsoids with signed distance fields (SDFs), achieving high-quality geometric reconstructions and rendering results.
+Other approaches have leveraged ray-Gaussian intersections for improved accuracy.
+Rade-GS [51], GOF [50], and PGSR [5] computed ray-ellipsoid intersections to obtain unbiased depth estimates and applied normal consistency supervision from 2DGS [16] to achieve state-of-the-art reconstruction.
+2DGS further refined this approach by flattening Gaussian ellipsoids into disks, better aligning primitives with surfaces while introducing additional regularization losses for enhanced geometric constraints.
+MVG-splatting [25] extended 2DGS by incorporating multi-view consistency constraints, enabling more complex surface reconstructions.
+In this work, we propose Quadratic Gaussian Splatting (QGS), a novel surface representation based on quadric surfaces designed to improve the geometric fitting of primitives.
+We establish Gaussian distributions on quadrics, enabling end-to-end optimization.
+Additionally, by integrating multi-view consistency supervision, QGS achieves state-of-the-art geometric reconstruction and high-quality rendering.
+3 Method
+3.1 Preliminary
+Kerbl et al. [21] proposed representing a scene using 3D Gaussian ellipsoids as primitives and render images using differentiable volume splatting.
+$C(\mathbf{p})=\sum_{i=0}^{N-1}G_{i}^{2D}(\mathbf{p})\alpha_{i}c_{i}\prod_{j=0}^{i-1}(1-G_{j}^{2D}(\mathbf{p})\alpha_{j})$ (1)
+Here, $\alpha_{i}$ represents the opacity, $c_{i}$ denotes the color of each Gaussian primitive modeled by spherical harmonics, which is also used in our method, and $\mathbf{p}$ is the pixel coordinate.
+Finally, each Gaussian primitive is optimized by minimizing the photometric loss.
+GS mesh reconstruction.
+Vanilla 3DGS [21] can render high-quality images, but it yields suboptimal results for scene geometry reconstruction.
+In subsequent surface reconstruction methods, volumetric approaches like GOF [50] and RadeGS [51] employ ray-splat intersection techniques, achieving SOTA reconstruction quality but limiting the consistency of normal and depth.
+In contrast, 2DGS [16] defines the 2D Gaussian distribution within a planar disk, which inherently provides consistent normals and depths across multiple views.
+But the disk is only a first-order approximation of the surface, which often leads to overly smooth reconstruction results in 2DGS, as shown in Fig 2.
+Figure 3: Illustration of a Gaussian on a quadric surface. The blue dot marks the Gaussian centroid, the red dot indicates the ray-splat intersection, and the red line represents the geodesic.
 3.2 Quadratic Gaussian Splatting
@@
 We will first introduce the Quadric Gaussian Model, then discuss the splatting design for quadrics, and finally explain the optimization process.
-3.2.1 Quadratic Gaussian Model
-Quadratic Model. Given a homogeneous coordinate $\mathbf{x}=[x,y,z,1]^{T}\in\mathbb{R}^{4}$, a quadric surface can be defined as the solution set to the following equation:
+Quadratic Gaussian Model. Given a homogeneous coordinate $\mathbf{x}=[x,y,z,1]^{T}\in\mathbb{R}^{4}$, a quadric surface can be defined as the solution set to the following equation:
 $f(x,y,z)$ $=Ax^{2}+2Bxy+2Cxz+2Dx+Ey^{2}$ (2)
@@
 $=\mathbf{x}^{T}\mathbf{Q}\mathbf{x}=0$
-Similar to [26], we apply congruent diagonalization to transform the above equation into its canonical form:
-$\mathbf{Q}=\mathbf{T}^{-T}\mathbf{D}\mathbf{T}^{-1},\mathrm{with}$ $\mathbf{D}\;\mathrm{diagonal},d_{ii}\in\{0,\pm 1\}$ (5)
-$\mathbf{T}=\left[\begin{matrix}\mathbf{u}&\mathbf{v}&\mathbf{w}&\mathbf{c}\\ 0&0&0&1\end{matrix}\right]$
-Here, $\mathbf{c}$ denotes the position of the quadric, and $[\mathbf{u},\mathbf{v},\mathbf{w}]$ can be decomposed into $\mathbf{RS}$, where $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}]$ and $\mathbf{S}=\mathrm{diag}(s_{1},s_{2},s_{3})$, which denotes the orientation and scale of the quadric in the object space.
-The matrix $\mathbf{D}$ defines the surface shape: $\mathbf{D}=\mathrm{diag}(1,1,1,1)$ yields an ellipsoid, while $\mathbf{D}=\mathrm{diag}(1,0,0,0)$ produces a plane.
+We apply a congruent transformation to obtain the transformation from object space to parameter space.
+$\mathbf{Q}=\mathbf{T}^{-T}\mathbf{D}\mathbf{T}^{-1},\mathrm{with}$ $\mathbf{D}\;\mathrm{symmetrical}$ (3)
+$\mathbf{T}=\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]$
+Here, $\mathbf{c}$ denotes the position of the quadric, and $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}]$ denotes the orientation of the quadric in the object space.
+The matrix $\mathbf{D}$ defines the surface scale and shape: $\mathbf{D}=\mathrm{diag}(s_{1}^{2},s_{2}^{2},s_{3}^{2},1)$ yields an ellipsoid, while $\mathbf{D}=\mathrm{diag}(s_{1}^{2},s_{2}^{2},0,0)$ produces a plane.
 To compute the Gaussian weight at any surface point, we first define a measure on the surface to establish the Gaussian distribution.
-To concentrate the Gaussian energy on the surface, we use geodesic length [27] as the metric: the geodesic distance between two surface points is the shortest path along the surface, as illustrated by the red line in Fig 3.
+A straightforward approach is to construct the Gaussian distribution using Euclidean distance.
+However, this leads to inconsistencies when the surface undergoes deformation, resulting in uneven splatting reconstruction artifacts, as shown in Fig.
+5.
+To address this, we use geodesic distance to ensure that weights adapt consistently to surface deformations.
+Fig.
+4 shows a 1D illustration.
+Figure 4: Density distribution among deformable primitives (disk→paraboloid).
+(a) Euclidean distance fails to adapt to curvature during primitive deformation, distorting density distributions (uneven weight in the blue point).
+(b) Our geodesic distance follows intrinsic curvature, maintaining spatial coherence across deformation and preventing uneven splatting artifacts.
 However, not all geodesics have closed-form solutions.
 When $\mathbf{Q}$ is an ellipsoid or hyperboloid, computing the geodesic length typically requires numerical methods.
-Therefore, we focus only on the case of a paraboloid:
-$f(x,y,z)$ $=\mathbf{x}^{T}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-T}\bar{D}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-1}\mathbf{x}$ (6)
+Luckily, as we aim at surface reconstruction, we focus on paraboloid with the following form:
+$f(x,y,z)$ $=\mathbf{x}^{T}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-T}\mathbf{D}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-1}\mathbf{x}$ (4)
 $=\hat{\mathbf{x}}^{T}\left[\begin{matrix}\frac{d_{11}}{s_{1}^{2}}&0&0&0\\ 0&\frac{d_{22}}{s_{2}^{2}}&0&0\\ 0&0&0&-\frac{d_{33}}{2s_{3}}\\ 0&0&-\frac{d_{33}}{2s_{3}}&0\end{matrix}\right]\hat{\mathbf{x}}$
@@
 To introduce a sign, we add another variable $t$ to control the sign, i.e., $s(x,t)=\tanh(t)\exp(x)$.
-Gaussian Distribution on Quadric.
-We will now describe how to define a Gaussian distribution on a paraboloid.
-First, we need to define a measure on the quadric.
-Paraboloid (Equation 7) could be expressed in explicit form:
+(a) w/ Euclid
+(b) w/ Geodesic
+Figure 5: Comparison between Euclidean and geodesic distances.
+The first row presents schematic illustrations, while the second row shows normal comparisons.
+The left column uses Euclidean distance, while the right column employs geodesic distance.
+Gaussian Distribution on Quadric. We will now describe how to define a Gaussian distribution on a paraboloid using geodesic lines. First, paraboloid (Equation 5) could be expressed in explicit form:
 $\hat{z}(\hat{x},\hat{y})=s_{3}(\frac{\mathrm{sign}(s_{1})}{s_{1}^{2}}\hat{x}^{2}+\frac{\mathrm{sign}(s_{2})}{s_{2}^{2}}\hat{y}^{2})$ (6)
 We convert it isometrically into cylindrical coordinates, i.e. $\hat{x}=\rho\cos\theta,\hat{y}=\rho\sin\theta$. And we rewrite Equation 6 as:
-$\vspace{-5pt}\begin{split}\hat{z}(\theta,\rho)=\;&s_{3}(\frac{\mathrm{sign}(s_{1})\cos^{2}\theta}{s_{1}^{2}}+\frac{\mathrm{sign}(s_{2})\sin^{2}\theta}{s_{2}^{2}})\rho^{2}\\ =\;&a(\theta)\rho^{2}\end{split}\vspace{-5pt}$ (9)
+$\begin{split}\hat{z}(\theta,\rho)=\;&s_{3}(\frac{\mathrm{sign}(s_{1})\cos^{2}\theta}{s_{1}^{2}}+\frac{\mathrm{sign}(s_{2})\sin^{2}\theta}{s_{2}^{2}})\rho^{2}\\ =\;&a(\theta)\rho^{2}\end{split}$ (7)
 Due to the paraboloid’s symmetry, for any point $\hat{\mathbf{p}}_{0}=(\rho_{0},\theta_{0},\hat{z}(\theta_{0},\rho_{0}))$, the intersection line of the plane $\theta=\theta_{0}$ with the paraboloid: $\hat{z}(\theta_{0},\rho),\rho\in(0,\rho_{0})$, is the geodesic to the origin.
 Then the geodesic distance is the arc length $l$ of this curve, as shown in Fig 3.
-Figure 3: Illustration of a Gaussian on a quadric surface. The blue dot marks the Gaussian centroid, the red dot indicates the ray-splat intersection, and the red line represents the geodesic.
 $l(a,\rho_{0})$ $=\int_{0}^{\rho_{0}}\sqrt{1+(2at)^{2}}\mathrm{d}t$ (8)
@@
 This indicates that 2DGS can be regarded as a specific degenerate case of QGS, whose more generalized nature allows it to fit high-curvature regions effectively.
+(a) normal
+(b) curvature
+Figure 6: The (a) normal map and (b) curvature map of a scene. In the curvature map, blue indicates higher curvature and curved areas, while red indicates lower curvature and flatter areas.
 3.2.1 Splatting
@@
 By substituting $\hat{\mathbf{p}}$ into the Equation 6, we solve a quadratic equation to obtain two intersection points: the nearer point $\hat{\mathbf{p}}_{n}=(\hat{x}_{n},\hat{y}_{n},\hat{z}_{n})$ and the farther point $\hat{\mathbf{p}}_{f}=(\hat{x}_{f},\hat{y}_{f},\hat{z}_{f})$, with $t_{n}\leq t_{f}$.
-To ensure multi-view consistency in QGS, we select one of these two points: First, if the geodesic distance of $\hat{\mathbf{p}}_{n}$ is within $3\sigma(\theta_{n})$, we choose $\hat{\mathbf{p}}_{n}$.
+To ensure multi-view consistency in QGS, only one valid point of the two is selected.
+First, if the geodesic distance of $\hat{\mathbf{p}}_{n}$ is within $3\sigma(\theta_{n})$, we choose $\hat{\mathbf{p}}_{n}$.
 If not, we check if $\hat{\mathbf{p}}_{f}$ is within $3\sigma(\theta_{f})$.
@@
 If neither condition is met, we assume no intersection between the ray $\hat{\mathbf{p}}(t)$ and the primitive.
+This assumption is generally valid because the significant weight typically causes the first point to occlude the second point.
 The derivation is provided in the supplementary material.
@@
 $\hat{K}_{0}(\hat{\mathbf{p}}_{0})=\frac{4\lambda_{1}\lambda_{2}}{(1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2})^{2}}$ (13)
-We render the normal map $N(\mathbf{p})$ and curvature map $K(\mathbf{p})$ by Equation 16 for a given viewpoint using alpha-blending, as shown in the Fig 4.
-$N(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}\textbf{n}_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$ (16)
-$K(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}K_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$
-normal
-curvature
-Figure 4: The normal map (left) and curvature map (right) of a scene. In the curvature map, blue indicates higher curvature and curved areas, while red indicates lower curvature and flatter areas.
-Per-tile Sorting and Per-pixel Resorting.
-In Equation 16, the index $i$ is determined by sorting primitives from near to far.
-Most GS-like methods sort by splat centroid depth rather than ray-splat intersection depth, which, as noted by StopThePop [25], causes popping artifacts.
-Additionally, we observed that centroid sorting introduces streak artifacts in geometry reconstruction, as shown in Fig 8.
-To address this, we apply StopThePop’s Per-tile Sorting and Per-pixel Resorting [25] to our QGS.
-In short, we first compute the ray closest to the quadric center in each tile to estimate intersection depth for rough sorting.
-During volume rendering, we use a buffer array to locally resort quadrics based on the intersection depth of each pixel.
-Further details are provided in the supplementary material.
+We render the normal map $\mathbf{N}$ and curvature map $\mathbf{K}$ by Equation 14 for a given viewpoint using alpha-blending, as shown in the Fig 6.
+$\{\mathbf{N},\mathbf{K}\}$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})\{\textbf{n}_{i},K_{i}\}$ (14)
+(a) w/o resorting
+(b) w/ resorting
+Figure 7: Comparison of depth normals with and without resorting. Centroid-only sorting (a) introduces stripe artifacts, while resorting (b) produces smoother and more consistent normals.
+Per-pixel Resorting.
+We adopt the per-pixel resorting method from Stopthepop to address the depth pop-out issue, as illustrated in Fig 7.
+Implementation and derivation details are provided in the supplementary material.
 3.3 Optimization
-For QGS, which has stronger geometric fitting capability, relying solely on photometric consistency error can introduce geometric noise, such as unsmooth surfaces.
-Therefore, we adopt the depth distortion loss and normal consistency loss proposed by 2DGS [12] to mitigate this.
-Depth Distortion. We employ the depth distortion loss to promote tighter alignment of the quadric surfaces.
+We now introduce the our training objectives during optimization.
+Depth Distortion. We adopt the depth distortion loss and normal consistency loss proposed by 2DGS [16].
 $\mathcal{L}_{d}=\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}\omega_{i}\omega_{j}(t_{i}-t_{j})^{2}$ (15)
 Here, $\omega_{i}=\bar{\alpha}_{i}T_{i}$ denotes the alpha-blending weight of the i-th Gaussian, and $t_{i}$ represents the depth at the ray-splat intersection.
-$i,j$ index the Gaussians contributing to the ray.
-Additionally, following GOF’s recommendation, we freeze the gradient of the depth distortion loss through $\omega_{i}$, optimizing only the depth $t_{i}$.
-Normal Consistency. 2DGS [12] introduces the normal consistency loss to ensure that all primitives on a ray are locally aligned with the actual surface.
+$i,j$ index the Gaussians along the ray.
+Curvature Guided Normal Consistency. 2DGS [16] introduces the normal consistency loss to ensure that all primitives on a ray are locally aligned with the actual surface.
 $\mathcal{L}_{n}$ $=\sum_{i}\omega_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$ (16)
 Here, $\mathbf{n}_{i}$ represents the splat normal facing the camera, while $\mathbf{N}$ is computed by differentiating the depth point $\mathbf{p}$ from neighboring pixels.
-$\mathbf{N}(u,v)=\frac{\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}}{|\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}|}$ (19)
-However, neighboring pixels may not satisfy the local planar assumption, especially in regions with significant depth variation.
-Using differential normal consistency loss in edge areas can introduce errors, contributing to the over-smoothing observed in 2DGS.
-PGSR [4] and MVG-splatting [19] have both noted this issue and use image edges to approximate geometric edges, applying additional processing at these edges.
-We have found that, in most scenes, image edges do not fully correspond to geometric edges, especially in uniformly lit areas.
-Thus, we utilize the curvature map, which more accurately corresponds to geometric edges and is both efficiently and uniquely generated by QGS, to guide normal supervision.
+However, neighboring pixels may violate the local planar assumption, particularly in edge regions.
+Using normal consistency loss in edge areas can introduce errors, contributing to the over-smoothing observed in 2DGS.
+A straightforward approach is to approximate geometric edges using image edges, which can then guide normal consistency supervision [5, 25].
+However, we found that, in most scenes, image edges do not fully correspond to geometric edges, especially in uniformly lit areas.
+Thus, we use the curvature map, which more accurately corresponds to geometric edges and is efficiently and uniquely generated by QGS, to guide normal supervision.
 $\lambda_{K}(K(u,v))=$ $1-\mathrm{sigmoid}(\ln(|K(u,v)|+\varepsilon)$ (17)
 $\mathcal{L}_{Kn}(u,v)$ $=\lambda_{K}(K(u,v))\mathcal{L}_{n}(u,v)$
+Multi-view regularization Loss.
+To ensure a fairer comparison with plane-based PGSR [5], we incorporate PGSR’s multi-view regularization, computing warp and photometric consistency losses via homography transformations.
+$\mathbf{H}_{rn}=$ $\mathbf{K}_{n}(\mathbf{R}_{rn}+\frac{\mathbf{T}_{rn}\mathbf{n}_{r}^{T}}{\mathbf{n}_{r}^{T}\mathbf{X}_{r}})\mathbf{K}_{r}^{-1}$ (18)
+[…]
+$(1-NCC(\mathbf{I}_{r}(\mathbf{p}_{r}),\mathbf{I}_{n}(\mathbf{H}_{rn}\mathbf{p}_{r})))$
+Here, $\mathbf{n}_{r}$ represents the normal map, and $\mathbf{X}_{r}$ denotes the 3D points projected from the depth map.
+$\mathbf{R}_{rn}$ represents the relative rotation, $\mathbf{K}$ is the camera intrinsic, $\mathbf{p}_{r}$ denotes the pixel coordinates, and $\mathbf{I}$ refers to the image intensity.
+This loss enforces multi-view constraints by computing the reprojection error from the depth and normal maps, along with normalized cross-correlation (NCC) [47].
 Final Loss. Finally, we input a sparse point cloud and posed images to optimize QGS with the following loss function:
-$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{Kn}$ (21)
+$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{Kn}+\lambda_{Mv}\mathcal{L}_{Mv}$ (19)
 4 Experiment
-Figure 5: Qualitative geometric reconstruction comparisons on three datasets (DTU, Tanks and Temples, and Mip-NeRF 360) against 2DGS and GOF, covering indoor, outdoor, object-centric, and forward-facing scenes.
-QGS shows obvious improvements in both detail and smoothness compared to 2DGS and GOF.
-CD (mm)$\downarrow$ 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean Time
-NeuS [30] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 $>$12h
-VolSDF [32] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 $>$12h
-Neuralangelo [18] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 $>$128h
+scan24
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+scan65
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+scan118
+Reference Image
+Refer to captionRefer to caption
+2DGS
+Refer to captionRefer to caption
+GOF
+Refer to captionRefer to caption
+PGSR
+Refer to captionRefer to caption
+Ours
+Figure 8: Qualitative geometric reconstruction comparisons on the DTU dataset. Our method achieves reconstructions of higher quality and greater detail.
+[…]
+NeuS [42] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 >>12h
+VolSDF [45] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 >>12h
+Neuralangelo [24] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 >>128h
+3DGS [21] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 11.2min
+Gaussian surfels [8] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 6.7min
 SuGaR[13] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 1h
-2DGS [12] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 0.32h
+2DGS [16] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 19.2min
 GOF [50] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 1h
-Ours 0.46 0.81 0.40 0.42 1.07 0.80 0.79 1.25 0.95 0.67 0.62 1.27 0.40 0.60 0.49 0.73 0.83h
-Table 1: Quantitative Chamfer Distance comparison of QGS with GS-like and NeRF-like methods on the DTU dataset.
-Compared to the planar representation in 2DGS, the surface-based QGS shows a clear improvement in geometric reconstruction quality.
-Additionally, QGS is an order of magnitude faster than NeRF-like methods.
-F1-Score $\uparrow$ NeuS Geo-NeuS Neuralangelo 2DGS GOF Ours
-Barn 0.29 0.33 0.70 0.41 0.51 0.43
-Caterpillar 0.29 0.26 0.36 0.24 0.41 0.31
-Courthouse 0.17 0.12 0.28 0.16 0.28 0.26
-Ignatius 0.83 0.72 0.89 0.52 0.68 0.79
-Meetingroom 0.24 0.20 0.32 0.17 0.28 0.25
-Truck 0.45 0.45 0.48 0.45 0.58 0.60
-Mean 0.38 0.35 0.50 0.33 0.46 0.44
-Time $>$24h $>$24h $>$127h 0.57h 2h 2h
-Table 2: Quantitative F1-Score comparison of QGS with GS-like and NeRF-like methods on the TNT dataset. QGS shows notable gains over 2DGS and is competitive with prior SOTA methods.
+GSDF [48] 0.59 0.94 0.46 0.38 1.30 0.77 0.73 1.59 1.29 0.76 0.59 1.22 0.38 0.52 0.51 0.80 32min
+GS-pull [52] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75 22min
+Ours w/o MV 0.46 0.76 0.40 0.38 0.92 0.80 0.76 1.25 0.95 0.67 0.62 1.20 0.38 0.60 0.47 0.71 25min
+PGSR [5] 0.40 0.60 0.39 0.37 0.78 0.59 0.53 1.18 0.67 0.63 0.48 0.62 0.34 0.42 0.39 0.56 40min
+Ours 0.38 0.62 0.37 0.38 0.75 0.55 0.51 1.12 0.68 0.61 0.46 0.58 0.35 0.41 0.40 0.54 48min
+Table 1: Quantitative Chamfer Distance comparison on the DTU dataset.
+Our method achieves the best performance both with and without multi-view regularization.
+,  ,   denote the best, second best, and third best results, respectively.
+F1-Score $\uparrow$ Geo-NeuS N-angelo 2DGS GOF Ours w/o MV PGSR Ours
+Barn 0.33 0.70 0.41 0.51 0.46 0.52 0.55
+Caterpillar 0.26 0.36 0.24 0.41 0.32 0.38 0.40
+Courthouse 0.12 0.28 0.16 0.28 0.26 0.26 0.28
+Ignatius 0.72 0.89 0.52 0.68 0.79 0.77 0.81
+Meetingroom 0.20 0.32 0.17 0.28 0.25 0.29 0.31
+Truck 0.45 0.48 0.45 0.58 0.60 0.62 0.64
+Mean 0.35 0.50 0.33 0.46 0.45 0.47 0.50
+Time >>24h >>127h 34min 114min 43min 66min 75min
+Table 2: Quantitative F1-Score comparison of QGS with GS-like and NeRF-like methods on the TNT dataset. QGS outperforms all methods, achieving state-of-the-art reconstruction results.
 We compared our method with several SOTA approaches across multiple datasets, including DTU [19], TNT [22], and Mip-NeRF 360 [2].
@@
 4.1 Implementation Details.
-Rasterizer We implemented Quadratic Splatting with custom CUDA kernels on the 3DGS framework, extending the renderer to output depth distortion maps, depth maps, normal maps, and curvature maps.
-Given the non-convex nature of paraboloids and the lack of an inverse for the geodesic distance function, we used rectangular truncation and approximations to compute image bounding boxes, as detailed in the supplementary materials.
-To handle the more complex surface intersections, we incorporate Per-tile Sorting, Per-pixel Local Resorting from StopThePop [25], with further details also in the supplementary materials.
+Rasterizer.
+We implemented Quadratic Splatting with custom CUDA kernels on the 3DGS framework [21], extending the renderer to output depth distortion maps, depth maps, normal maps, and curvature maps.
+Since quadrics can be non-convex, we used rectangular truncation and approximations to compute image bounding boxes, as detailed in the supplementary materials.
 Settings.
 We adopted the adaptive control strategy from 3DGS [21].
-Similar to 2DGS [12], the QGS splatting process approximates by projecting the 3D center gradient onto screen space instead of relying directly on the 2D projected center gradient, using a gradient threshold of 0.3 and a percent dense value of 0.001.
-To ensure a similar number of Gaussian primitives across methods, we also set the gradient threshold for other methods to 2e-4.
+Similar to 2DGS [16], QGS projects the 3D center gradient onto screen space instead of using the 2D projected gradient.
+A gradient threshold of 0.3 and a percent dense value of 0.001 ensure consistent point cloud number with other methods.
 All experiments were conducted on a single A6000 GPU.
 Mesh Extraction.
-We used a weighted combination of 0.5 median depth (i.e., $t_{\mathrm{median}}=\max{t_{i}|T_{i}>0.5}$) and 0.5 expected depth (i.e., $t_{\mathrm{expected}}=\sum_{i=0}^{N-1}G_{i}\alpha_{i}T_{i}t_{i}$) as the final output depth.
+We used median depth (i.e., $t_{\mathrm{median}}=\max{t_{i}|T_{i}>0.5}$) as the final output depth.
 Then we fused the depth maps using Truncated Signed Distance Fusion (TSDF) with Open3D [53].
-Following the 2DGS setup, we set the voxel size to 0.004 and truncation threshold to 0.002 for fair comparison.
-4.2 Comparison
-Geometry Evaluation In all experiments, we performed evaluations at half image resolution.
-In Table 1, we compared QGS with implicit [30, 32, 18] and explicit [12, 35, 11] SOTA reconstruction methods on the DTU dataset using Chamfer Distance as the benchmark.
-QGS outperformed all explicit methods, and while a gap remains with Neuralangelo [18], our reconstruction time is two orders of magnitude faster.
-In certain scenes, such as Scan 97 and Scan 105, QGS even outperformed Neuralangelo.
-In Table 2, we evaluated QGS and other SOTA methods on the TNT dataset using F1-score.
-In large-scale scenes, QGS consistently outperformed 2DGS [12].
-Notably, in the Courthouse and Meetingroom scenes, QGS captured more details due to its higher geometric flexibility, as shown in the Fig 6.
-2DGS
-QGS
-Figure 6: Mesh reconstruction comparison between QGS and 2DGS in large-scale scenes.
-The first row shows the Courthouse scene, with both using median depth, and the second row shows the Meetingroom scene, using a 0.5 weighted combination of median and expected depth.
-As shown in Fig 5, our QGS achieves more complete geometry and stronger detail capture due to the greater flexibility of individual primitives.
-Note that our mesh extraction is limited by the lower resolution of the voxel grid in TSDF fusion, resulting in slightly lower accuracy compared to GOF [35], which uses marching tetrahedra, leading to longer meshing times and higher VRAM consumption.
-QGS’s second-order fit greatly improves on 2DGS [12], capturing finer details and providing second-order geometric information.
-The curvature map in Fig 4 illustrates how QGS better conforms to object edges with higher curvature, while adhering to flatter surfaces in other areas.
-It is important to note that the advantage of surface-based representations lies in their ability to provide multi-view consistent depth and normals (and curvature in our method).
-Thus, both theoretically and practically, it is straightforward to transfer multi-view geometric techniques, as done in PGSR [4] and MVG-Splatting [19], to QGS.
-We leave this exploration for future work to further enhance our approach.
-Rendering Evaluation We compared rendering quality on the Mip-NeRF360 dataset with baseline approaches.
+Following the 2DGS setup, we set the voxel size to 0.004 and truncation threshold to 0.02.
+4.2 Comparison.
+Geometry Evaluation.
+In all experiments, evaluations were performed at half the image resolution.
+To clearly demonstrate the improvement introduced by our quadric surfels and to maintain consistency with other methods, we denote the variant of QGS without multi-view regularization loss as QGS w/o MV.
+In Table 1, we compared QGS with implicit [42, 45, 24] and explicit [16, 50, 13, 48, 5, 52] SOTA reconstruction methods on the DTU dataset using Chamfer Distance as the metric.
+As shown in Table 1, our method outperforms both implicit and explicit approaches on the DTU dataset.
+Furthermore, even without multi-view regularization, our method surpasses others in accuracy while maintaining competitive speed.
+This improvement stems from the superior geometric fitting capability of quadrics, which enables finer detail preservation, as illustrated in Fig 8.
+In Table 2, we also evaluated QGS and other SOTA methods [11, 24, 16, 50, 5] on the TNT dataset using F1-score.
+Without multi-view regularization, our method significantly outperforms the disk-based 2DGS and achieves comparable results to GOF while requiring only half the computation time.
+With multi-view regularization, it surpasses all methods in reconstruction accuracy.
+Additionally, we provide a qualitative comparison between QGS without multi-view regularization and 2DGS to further demonstrate the enhanced geometric fitting capability of quadric surfels, as shown in Fig 9.
+(a) 2DGS
+(b) QGS w/o MV
+Figure 9: Mesh reconstruction comparison between 2DGS and QGS w/o MV. The first row shows the Meetingroom in TNT dataset, and the second row shows the kitchen scene in Mip-NeRF 360 dataset.
+Rendering Evaluation.
+We compared rendering quality on the Mip-NeRF360 dataset with baseline approaches.
 As a surface-based method, QGS achieves competitive rendering results in indoor scenes, as shown in Table 3, but performs less effectively in outdoor environments.
@@
 Mip-NeRF360 31.72 0.917 0.180 24.47 0.691 0.283
-3DGS 30.99 0.926 0.199 24.24 0.705 0.283
+3DGS 30.52 0.921 0.199 24.45 0.728 0.240
 SuGar 29.44 0.911 0.216 22.76 0.631 0.349
@@
 GOF 30.80 0.928 0.167 24.76 0.742 0.225
-QGS 30.76 0.926 0.166 24.08 0.699 0.273
+Ours w/o MV 30.48 0.926 0.166 24.56 0.724 0.239
+PGSR 30.35 0.924 0.176 24.29 0.718 0.236
+Ours 30.45 0.919 0.184 24.32 0.706 0.242
 Table 3: Quantitative comparison of appearance between QGS, GS-like, and NeRF-like methods on the Mip-NeRF 360 dataset.
 4.3 Ablation
-Curvature-Guided Normal Consistency Loss.
-As mentioned in Section 3.2, we use curvature to guide normal consistency supervision.
-High-curvature regions are treated as high-frequency areas that violate the local planar assumption, so we reduce supervision in these regions.
-As shown in Table 4 and Fig 7, curvature-guided normal supervision helps recover more scene details, leading to improved geometric quality.
-w/o $\lambda_{K}$
-w/ $\lambda_{K}$
-Figure 7: Comparison of curvature-guided normal consistency versus without curvature guidance. Curvature-guided supervision effectively reduces over-smoothing and recovers more details.
-F1-Score $\uparrow$ Truck Ignatius
+In this section, we assess the impact of individual QGS components on reconstruction quality, including per-pixel resorting, curvature-guided normal consistency, and multi-view regularization, as presented in Table 4 and Fig 10.
+We denote w/o sort as the absence of per-pixel resorting, w/o $\lambda_{K}$ as the absence of curvature-guided normal consistency, and w/o MV as the absence of multi-view regularization.
+We observe that: (a) Disabling curvature guidance fills fine gaps in the normal map, reducing reconstruction quality.
+(b) The absence of per-pixel sorting similarly fills gaps and leads to uneven surfaces.
+(c) Disabling multi-view regularization causes overfitting in low-texture regions, resulting in surface artifacts, such as dents on the front of the truck.
+Additionally, we conducted ablation studies without multi-view regularization to isolate and evaluate the effectiveness of quadrics.
+Specifically, we fixed the quadric’s z-axis scale to 0.001, approximating it as a disk (denoted as $s_{3}$ fixed).
+We also evaluated optimization using Euclidean distance (denoted as w/ Euclid).
+Both modifications significantly reduced reconstruction quality, as shown in Table 5.
+F1-Score $\uparrow$ B C CH I M T mean
+Full model 0.55 0.40 0.28 0.81 0.31 0.64 0.50
+w/o Sort 0.53 0.39 0.26 0.77 0.29 0.62 0.48
+w/o $\lambda_{K}$ 0.56 0.37 0.27 0.80 0.29 0.62 0.49
+w/o MV 0.46 0.32 0.26 0.79 0.25 0.60 0.45
+Table 4: Quantitative ablation study on the TNT dataset.
+We present ablation results for Barn (B), Caterpillar (C), Courthouse (CH), Ignatius (I), Meeting Room (M), and Truck (T).
+Our full setting achieves the best performance.
+(a) w/o $\lambda_{K}$
+(b) w/o Sort
+(c) w/o MV
+(d) Full
+Figure 10: Ablation study. The full setting achieves the highest reconstruction quality.
+Setting w/o MV $s_{3}$ fixed w/ Euclid
 […]
-w/o $\lambda_{K}$ 0.63 0.51 0.57 0.74 0.75 0.75
-w/ $\lambda_{K}$ 0.65 0.55 0.60 0.78 0.80 0.79
-Table 4: Quantitative F1-Score comparison on the TNT dataset with and without curvature-guided normal supervision.
-Per-tile Sorting.
-2DGS suggests the actual surface lies at the median intersection point where opacity reaches 0.5.
-However, most GS-like methods sort Gaussians by centroid depth, leading to streak-like inconsistencies, as shown in Fig 8.
-To address this, we implemented per-tile sorting (TS) and per-pixel resorting (PR) based on the methods proposed in StopThePop [25], further refining them to remove inconsistencies and enhance reconstruction quality, as shown in Table 5.
-w/o TS and PR
-w/ TS
-w/ TS and PR
-Figure 8: Normal comparison with and without per-tile sorting. Centroid depth sorting causes noticeable streak-like inconsistencies, also seen in 2DGS.
-CD (mm)$\downarrow$ 24 37 97 122
-w/o TS and PR 0.48 0.85 1.04 0.55
-w/ TS 0.46 0.80 0.96 0.52
-w/ TS + PR 0.46 0.81 0.95 0.49
-Table 5: Quantitative Chamfer Distance comparison on the DTU dataset with and without per-tile sorting.
+Table 5: Ablation study without multi-view regularization on the TNT dataset. Fixing the quadric as a disk or using Euclidean distance optimization both result in reduced reconstruction quality.
 5 Conclusion
@@
 Additional results are in the supplementary materials.
-Discussion.
-Gaussian Splatting is a rapidly evolving field, with numerous works exploring its use for surface reconstruction, all developed independently from ours, such as Rade-GS [36], PGSR [4], Trim-GS [7], and MVG-Splatting [19].
-Our approach differs in focus, introducing a novel representation that provides curvature information.
-Our contributions are orthogonal to these approaches and, in theory, could enhance their methods as well.
-Limitation.
-Since the bounding box of QGS primitives on the image is approximated, QGS processes more tiles than typical GS methods, particularly for larger quadric surfaces, resulting in a speed disadvantage.
-As a surface-based representation, QGS is still limited in capturing foggy scenes or transparent objects.
+Acknowledgement
+This work was supported by the National Natural Science Foundation of China (Grants U22B2055 and 62273345), the Beijing Natural Science Foundation (Grant L223003), and the Key R&D Project in Henan Province (Grant 231111210300).
 Supplementary Material
-In this supplementary material, we provide detailed explanations of the following: (1) Calculation of the projected bounding box of primitives.
+In this supplementary material, we provide detailed explanations of the following: (1) Comparison of quadric (QGS) and disk (2DGS [16]).
+(2) Calculation of the projected bounding box of primitives.
 (3) Details of per-tile and per-pixel sorting.
@@
 (7) We present more qualitative results.
-Additionally, we include a video (QGS.mp4) that briefly explains the main workflow of QGS and showcases comparative results.
+Comparison of Disk and Quadric
+We visualize the primitives of QGS (Quadric) and 2DGS (Disk), as shown in Fig 11.
+It can be observed that quadrics tightly conform to the surface in a curved manner, with a more noticeable fit in regions of high curvature.
+(a) Mesh
+(b) Disk
+(c) Quadric
+Figure 11: Visualizaion results on GSO dataset [9]. (a) shows the ground truth mesh, (b) displays the disk primitives, and (c) presents the quadric primitives.
 Details of Bounding Box Calculation.
 In 2DGS/3DGS [16, 21], Gaussian primitives are enclosed convex representations, making it straightforward and convenient to compute their bounding box on the image, as shown in the first column of Fig 12.
-Figure 9: Comparison diagram of the bounding boxes between 2DGS and QGS.
+Figure 12: Comparison diagram of the bounding boxes between loose and tight case.
 However, QGS includes concave cases, and geodesic distance complicates analytical determination of the Gaussian primitive’s rendered portion during preprocessing.
 Specifically, the geodesic function Equation 10 in the main text lacks an elementary inverse, so computing the equipotential $\rho(\theta_{0})=\{l^{-1}(l_{0}):l_{0}=\sigma(\theta_{0})\}$, requires numerical or approximate solutions.
-Even with an analytical solution for ${\rho(\theta_{0})=l^{-1}(\sigma(\theta_{0}))}$, the Gaussian distribution in non-Euclidean space means equipotential lines may not fully enclose the primitive, requiring extra boundary calculations.
-To simplify, we use the quadric’s circumscribing rectangular box as the 3D bounding box.
-Its 8 vertices are projected onto the image plane, and their 2D AABB is used as the 2D bounding box, as shown in Fig 9, columns two and three.
+Moreover, even with an analytical solution for $\rho(\theta_{0})=l^{-1}(\sigma(\theta_{0}))$}, the Gaussian distribution in non-Euclidean space means equipotential lines may not fully enclose the primitive, requiring extra boundary calculations.
+A straightforward approach is to use the projection of the quadric’s 3D bounding box as the 2D bounding box, as shown in Fig 12 (a).
+Its eight vertices are projected onto the image plane, and their 2D axis-aligned bounding box (AABB) is used as the 2D bounding box, as illustrated in Fig 12 (c).
+However, this loose case significantly enlarges the 2D bounding box, leading to a decrease in efficiency.
+To address this, we construct a tighter bounding box using the quadric’s tangent planes, as shown in Fig 12 (b).
 Specifically, we first scale the geodesic function to a quadratic polynomial function:
@@
 Figure 13: Comparison diagram of the approximate function versus the geodesic function. The red line represents the approximate function, while the black line represents the geodesic function.
-Since the quadratic curve passes through the origin, we take its positive root, representing the horizontal distance from the primitive’s vertex at a geodesic distance of $\sigma_{0}$.
+Since the Equation 20 through the origin, we take its positive root, representing the horizontal distance from the primitive’s vertex at a geodesic distance of $\sigma_{0}(\theta_{0})$.
 We compute this for the major and minor axes, i.e.
-$\theta=0$ and $\theta=\pi/2$, yielding $\rho_{l}$ and $\rho_{s}$, with circumscribing box radii $(\rho_{l},\rho_{s},a\max(\rho_{l}^{2},\rho_{s}^{2}))$.
-We then project the eight vertices of the bounding box onto the image and take the minimum 2D bounding box of the resulting polygon.
-Due to approximations, gaps may appear between the bounding box and the primitive as shown in the third column of Fig 9, affecting rendering speed but not quality.
+$\theta=0$ and $\theta=\pi/2$, yielding $\rho_{1}$ and $\rho_{2}$, and obtain four points: $(\pm\rho_{1},0,\rho_{1}^{2}),(0,\pm\rho_{2},\rho_{2}^{2})$.
+We then compute the intersection of the tangent planes at these four points with the plane $z=0$.
+$x_{1}=\pm\frac{\rho_{1}s_{1}-\rho_{1}^{2}}{s_{1}},\;\;x_{2}=\pm\frac{\rho_{2}s_{2}-\rho_{2}^{2}}{s_{2}}$ (22)
+At this stage, we obtain a 3D truncated pyramid with top dimensions $\rho_{1}\times\rho_{2}$, bottom dimensions $x_{1}\times x_{2}$, and height $\max(\rho_{1}^{2},\rho_{2}^{2})$.
+We then project its eight vertices onto the image plane and determine the minimum 2D bounding box enclosing the resulting polygon as shown in Fig 12 (d).
+Mip-NeRF360 TNT
+Storage FPS Training Time Storage FPS
+2DGS 556MB 15.34 1h5min 218MB 31.66
+GOF 825MB 7.61 2h12min 367MB 13.45
+QGS 688MB 7.92 1h48min 274MB 13.27
+QGS w/ TB 641MB 14.15 1h13min 263MB 25.36
+Table 6: Speed and storage comparison on the Mip-NeRF 360 [2] and TNT [22] datasets. With a tighter bounding box, QGS achieves a noticeable speed improvement.
+By introducing a tighter bounding box and optimizing global memory access on the GPU, QGS achieves twice the speed.
+We denote the accelerated QGS as QGS w/ TB and compare the storage and speed of different methods on the Mip-NeRF 360 and TNT datasets as shown in Table 6.
 Details of Per-tile Sorting and Per-pixel Resorting
@@
 However, for surface-based representations like 2DGS and QGS, the Gaussian distribution is concentrated on the surface, making the alpha-blending order more sensitive.
-Most reconstruction methods [35, 36, 19, 4], including 2DGS, sort by Gaussian centroid depth, causing geometric inconsistencies, as shown in Fig 8 of the main text.
 To address this, we introduce StopThePop’s per-tile sorting and per-pixel resorting [33] for more precise ordering.
@@
 Where $X_{i}$ represents properties such as the color, depth, normal, or curvature of the primitives. The gradient computation in a back-to-front order is given as follows:
-$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}$ $=X_{i}T_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{1-\bar{\alpha}_{i}}$
-$=(X_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
+$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}$ $=(X_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
 We rewrite it as front-to-back form:
@@
 $At^{2}+Bt+C=0$
-$t_{n}=\frac{-B-\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A},$
-$t_{f}=\frac{-B+\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A}$
+$t_{n,f}=\frac{-B\pm\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A}$
 However, Equation 24 encounters numerical issues when $|A|$ is very small, particularly during backpropagation, as a small denominator can cause floating-point overflow, resulting in invalid values such as NaN or Inf.
@@
 $\mathbf{\hat{n}}(\mathbf{p}_{o})=[\frac{2\mathrm{sign}(s_{1})\hat{o}_{1}}{s_{1}^{2}},\frac{2\mathrm{sign}(s_{2})\hat{o}_{2}}{s_{2}^{2}},-\frac{1}{s_{3}}]^{T}$
-The projection of $(\mathbf{p}_{o}-\hat{\mathbf{o}})$ onto the normal direction $\mathbf{n}$ is:
-$m=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{||\hat{\mathbf{n}}(\mathbf{p}_{o})||}$
+[…]
+[…]
 As indicated by the thick red line segment in the Fig 14. On the other hand, the cosine of the angle between the normal vector and the viewing direction can be expressed as:
-$\mathrm{cos}\theta=\frac{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}{||\mathbf{\hat{r}}||\cdot||\mathbf{\hat{n}}(\mathbf{p}_{o})||}$
-Noting that $||\hat{\mathbf{r}}||=1$, the length of the line from the camera center along the viewing direction to the intersection with the tangent plane is given by:
-$n$ $=\frac{m}{\cos\theta}=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}$
-$=-\frac{C}{B}$
+[…]
+Noting that $\hat{\mathbf{r}}||=1$, the length of the line from the camera center along the viewing direction to the intersection with the tangent plane is given by:
+$n$ $=\frac{m}{\cos\theta}=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}=-\frac{C}{B}$
 It is evident that when the surface is nearly flat or the viewing direction is almost perpendicular to the horizontal plane, the approximate solution nearly coincides with the exact solution.
@@
 For the geodesic arc length formula:
-[…]
+$l(a,\rho_{0})$ $=\int\sqrt{1+(2at)^{2}}dt$ (25)
 Let $u=2at,x=\arctan(u)$, then Equation 25 becomes:
-[…]
+$l(a,\rho_{0})$ $=\int\frac{\sqrt{1+u^{2}}}{2a}du$ (26)
 $=\frac{1}{2a}\int\sqrt{1+\tan^{2}(x)}d\tan(x)$
-$=\frac{1}{2a}\int\sec(x)d\tan(x)$
-[…]
+$=\frac{1}{2a}\int\sec^{3}(x)dx$
 Equation 26 can be solved using integration by parts:
-[…]
-[…]
-[…]
-[…]
-[…]
+$\int\sec^{3}(x)dx$ $=\tan(x)\sec(x)-\int\sec^{3}(x)dx$
+$+\int\sec(x)dx$
+Using $\int\sec(x)dx=\ln|\sec(x)+\tan(x)|+C$, we have:
+$\int\sec^{3}(x)dx$ $=\frac{\tan(x)\sec(x)+\ln|\sec(x)+\tan(x)|}{2}$
 $=(u\sqrt{1+u^{2}}+\ln(u+\sqrt{1+u^{2}}))/2$
@@
 The partial derivatives at $\hat{\mathbf{p}}_{0}$ are:
-$x_{u}$ $=(1,0,2\lambda_{1}\hat{x}_{0})$
-$x_{v}$ $=(0,1,2\lambda_{2}\hat{y}_{0})$
+$x_{u}=(1,0,2\lambda_{1}\hat{x}_{0}),\;\;x_{v}=(0,1,2\lambda_{2}\hat{y}_{0})$
 The first fundamental form is:
@@
 The second fundamental form is:
-$n=\frac{x_{u}\times x_{v}}{||x_{u}\times x_{v}||}$ $=\frac{(-2\lambda_{1}\hat{x}_{0},-2\lambda_{2}\hat{y}_{0},1)}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
-$x_{uu}$ $=(0,0,2\lambda_{1})$
-$x_{uv}$ $=(0,0,0)$
-$x_{vv}$ $=(0,0,2\lambda_{2})$
+[…]
+$x_{uu}$ $=(0,0,2\lambda_{1}),\;x_{uv}=(0,0,0),\;x_{vv}=(0,0,2\lambda_{2})$
 $L$ $=<n,x_{uu}>=\frac{2\lambda_{1}}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
@@
 In this section, we present additional qualitative results of QGS in both indoor and outdoor scenarios.
-Fig 12 compares QGS and 2DGS on the DTU dataset, where QGS demonstrates sharper contours.
+Figure 15 presents a comparison of QGS with 2DGS [16], GOF [50], and PGSR [5] on the DTU dataset, illustrating that QGS captures more geometric details.
 QGS also achieves accurate reconstructions in indoor and outdoor scenes, as shown in Fig 16 with results from the TNT and Mip-NeRF 360 datasets.
@@
 As shown in Fig 18, QGS captures more building details than 2DGS [16] in aerial views.
-Additionally, most methods using the default 3DGS configuration failed to reconstruct the Courthouse scene and the urban scene, both containing nearly 1,000 images.
-For these two scenes, we simply doubled the training parameters, setting the total iterations to 60,000, densifying every 800 iterations from the 4,000th to the 48,000th iteration, and resetting Gaussian opacity every 6,000 iterations.
-For other scenes, we used the default 3DGS configuration.
-Figure 12: Qualitative geometry comparison for the DTU dataset [14].
+scan37
+scan55
+scan63
+scan69
+scan110
+Reference Image
+2DGS [16]
+GOF [50]
+PGSR [5]
+Ours
+Figure 15: Qualitative geometric reconstruction comparisons on the DTU dataset [19]. Our method achieves reconstructions of higher quality and greater detail.
 Figure 16: Qualitative geometry results for the Tanks and Temples dataset [22] and Mip-NeRF 360 dataset [2].
@@
 Figure 18: Qualitative geometry result for the urban scene captured from aerial view, involving over 1,000 images.
-In the subfigure, the left shows the results of 2DGS, while the right shows the results of QGS.
+In the subfigure, the left shows the results of 2DGS [16], while the right shows the results of QGS.